### v3

# Do MLLMs Capture How Interfaces Guide User Behavior? A Benchmark for Multimodal UI/UX Design Understanding

Abstract
User interface (UI) design goes beyond visuals, guiding user behavior and overall user experience (UX). Strategically crafted interfaces, for example, can boost sign-ups and drive business sales, underscoring the shift toward UI/UX as a unified design concept. While recent studies have explored UI quality evaluation using Multimodal Large Language Models (MLLMs), they largely focus on surface-level features, overlooking behavior-oriented aspects. To fill this gap, we introduce WiserUI-Bench  [^1] https://github.com/jeochris/wiserui-bench, a novel benchmark for assessing models’ multimodal understanding of UI/UX design. It includes 300 diverse real-world UI image pairs, each consisting of two design variants A/B-tested at scale by actual companies, where one was empirically validated to steer more user actions than the other. Each pair is accompanied one or more of 684 expert-curated rationales that capture key factors behind each winning design’s effectiveness, spanning diverse cognitive dimensions of UX. Our benchmark supports two core tasks: (1) selecting the more effective UI/UX design by predicting the A/B test verified winner and (2) assessing how well a model, given the winner, can explain its effectiveness in alignment with expert reasoning. Experiments across several MLLMs show that current models exhibit limited nuanced reasoning about UI/UX design and its behavioral impact. We believe our work will foster research in UI/UX understanding and enable broader applications such as behavior-aware interface optimization.
1 Introduction
User interface (UI) design is central to application development. While visual aesthetics lay its foundation, the fundamental goal is to steer user behavior during interaction, such as encouraging sign-ups or purchases, and thereby boost service revenue (Fogg 2002). This behavioral focus expands to UI/UX design, uniting visual form with user experience (UX) considerations (Norman 1988).
Figure 1: A real-world behavior-driven design decision example from our benchmark, grounded in actual A/B test results. Such cases reflect UI changes that steer user actions, making them effective testbeds for evaluating models’ visual reasoning about UI/UX design.
Practitioners validate such decisions through large-scale A/B tests that randomly assign users to alternative UI variants and measure which one more effectively drives the desired action (Kohavi et al. 2009). These tests often reveal that even subtle design changes can significantly lift conversions. For example, as shown in Figure 1, simply reducing the number of options lowered user friction and increased sign-ups. Given that predicting and explaining such outcomes is nontrivial even for humans, hinging on complex cognitive processes of user experience and implicit human preferences, assessing multimodal understanding of UI/UX design becomes a crucial challenge, demanding not only identifications of visual elements but also reasoning about how design influences user behavior.
Benchmark # Samples Unit Source Quality Source Type
Large-Scale User
Validation
Expert
Rationale
Evaluation Objective
Yang and Li (2024)
382 Single Legacy UI Mobile
Guideline Violation
Detection
JitterWeb
(Wu et al. 2024a)
2.3M Pair Synthetic Mobile + Web
Basic Visual Quality
BetterWeb
(Wu et al. 2024a)
892 Pair Synthetic Mobile
Basic Visual Quality
UICrit
(Duan et al. 2024)
983 Single Legacy UI Mobile
UI Design Critique
WiserUI-Bench
300 Pair
Actual
Production
Mobile + Web
(Real A/B Test)
UI/UX Design Understanding
(User Behavior-Oriented)
Table 1: Comparison between WiserUI-Bench and existing UI evaluation benchmarks.
Recent efforts have explored how to evaluate UI design quality with Multimodal Large Language Models (MLLMs); however, they all fall short in capturing the behavioral impact of UI/UX design, as summarized in Table 1. While some focus solely on basic visual, heuristic-based attributes (Yang and Li 2024; Wu et al. 2024a), others rely on handwritten critiques of individual UIs, lacking validation against real user behavior data and offering limited coverage of cognitive aspects (Duan et al. 2024).
To address these limitations, we introduce WiserUI-Bench (Pairwise UI/UX Design Understanding Benchmark), a novel benchmark for evaluating a model’s multimodal understanding of UI/UX design. WiserUI-Bench includes 300 real-world UI image pairs with winning variants verified by A/B tests from actual companies, spanning varied contexts and sourced from reliable platforms. Each pair is accompanied by one or more expert-curated rationales, 684 in total, written post hoc with knowledge of the winner to explain its effectiveness, covering key behavioral drivers. An example instance is illustrated in the upper part of Figure 2. Uniquely grounded in large-scale real user behavior data, our benchmark mirrors authentic usage and preference patterns, providing a reliable basis for evaluation. Each rationale is tagged with one of the three cognitive UX dimensions articulated by Norman (2007)—perception, what users notice; memory, what they remember; and action, how they act—and with the specific UX law it reflects. This enables fine-grained assessment of how well models grasp the underlying user experience behind each design choice. For example, the action-related rationale in Figure 2 corresponds to Fitts’s Law, which states that users act faster when targets are larger and closer (Fitts and Peterson 1964).
Our benchmark supports two core tasks, illustrated in the lower part of Figure 2: (1) selecting the more effective UI/UX design from a given pair by predicting the A/B test verified winner, and (2) assessing how well model-generated rationales, given the winner, align with expert reasoning. This dual-task setup captures the essential capabilities for UI/UX assessment, not only identifying which design better guides user behavior, but also explaining why reliably.
Experiments on a broad range of proprietary and open-source MLLMs on WiserUI-Bench show that current models struggle with behavior-centered visual reasoning. Models perform near-random on selection, and their post-hoc reasoning remains insufficient to reach expert level. We expect that WiserUI-Bench and our novel UI/UX-focused tasks will spur further research on this critical problem and potentially push toward behavior-aware UI generation, the direction that truly matters for achieving user-centered design.
In summary, our main contributions are as follows:
  1. 1.
    We introduce a novel and challenging task of multimodal UI/UX understanding that centers on the behavioral impact of design.
  2. 2.
    We release WiserUI-Bench, a benchmark of 300 real-world UI pairs uniquely constructed from verified results of actual large-scale A/B tests conducted by real companies, along with expert-curated key rationales.
  3. 3.
    Through comprehensive experiments and analyses across diverse MLLMs, we uncover their limited ability to reason about behavior-driven UI effectiveness, highlighting a critical gap in current visual reasoning.
2 Related Works
Visual Reasoning
Visual reasoning benchmarks for Multimodal Large Language Models (MLLMs) evaluate models’ ability to process visual information and reason at a cognitive level. While some target general tasks (Yue et al. 2024; Liu et al. 2024b; Fu et al. 2024; Li et al. 2024a), others focus on specific domains like mathematics or science (Hao et al. 2025; Lu et al. 2024; Xu et al. 2025). However, few benchmarks address reasoning about human preferences in visual contexts. Moreover, although some studies explore visual reasoning across multiple images (Cheng et al. 2025; Zhao et al. 2024), such cases remain rare. Our task incorporates both aspects, reasoning over paired UI images and aligning with behavior-driven preferences, as captured by real-world A/B test outcomes.
Figure 2: Overview of WiserUI-Bench and two main tasks. Each instance contains a UI image pair with verified A/B test winner and expert-curated key rationales explaining it. These span the three cognitive UX dimensions; the example includes all three, though most cover fewer. MLLMs are evaluated on (1) selecting the more effective UI/UX design by predicting the verified winner, and (2) explaining the effectiveness of a given winner, measured by whether the model covers each expert rationale.
UI Evaluation
Recent advances have driven progress in evaluating UI quality with MLLMs. Yang and Li (2024) proposes a benchmark for detecting guideline violations in individual screens, while Wu et al. (2024a) generates synthetic UI pairs via perturbations such as color noise or spacing changes. However, both are limited to heuristic analysis of surface-level appearance. Duan et al. (2024) evaluates single screens using expert-critique annotations, but the benchmark has key limitations: it lacks validation with real-world user behavior and relies on RICO (Deka et al. 2017), an outdated, mobile-only dataset that limits its relevance to today’s diverse, production-level designs. Prior work also suggests that pairwise evaluation yields more reliable signals, particularly for subjective tasks like ours that target behavior-driven preferences captured by real-world A/B tests (Abel et al. 2020; Bawabe et al. 2021). In this context, WiserUI-Bench offers a stronger foundation by grounding design effectiveness in large-scale user behavior data and requiring models to engage in deeper, nuanced reasoning over both visuals and user experience in a pairwise format.
Simulation of Human Behavior
Efforts to simulate human behavior with AI are ongoing. Park et al. (2023) places LLM-driven characters in a sandbox town, producing socially plausible interactions. Lu et al. (2025a) moves closer to real behavior by using actual platform logs to predict next user actions from prior context, yet remains text-only. In the UI domain, Lu et al. (2025b) and Wang et al. (2025a) deploy persona-based agents to browse interfaces and examine how designs may steer user behavior; the former checks realism only via qualitative interviews, while the latter compares against actual A/B tests but in a single scenario. By contrast, our work grounds evaluation on numerous A/B test results from real users and introduces a novel multimodal task: reason over UI variant pairs to predict user behavior, enabling direct measurement of behavioral fidelity.
3 WiserUI-Bench
We propose WiserUI-Bench, a new benchmark for evaluating multimodal understanding of UI/UX design, a key factor in crafting interfaces that influence user behavior.
3.1 Benchmark construction
Data Sources
To construct a benchmark that faithfully reflects real-world UI/UX design decisions, we collect UI images from industry-validated A/B testing data. Specifically, we aggregate data from widely recognized A/B testing and analysis platforms [^2] https://vwo.com/success-stories/, https://goodui.org/leaks/, https://abtest.design/, which showcase large-scale A/B tests conducted by companies across various industries and regions. These sources provide both UI variants and verified outcomes based on actual user behavior data. By grounding our benchmark in such real, rigorously tested cases, we ensure a credible basis for evaluating UI/UX design understanding in practical scenarios.
Data Curation Pipeline
As illustrated in Figure 3, our curation process involves multiple stages to ensure high-quality annotations. First, we extract raw UI images and accompanying textual descriptions from our selected platform sources. Next, we refine these images by eliminating any visual markers, such as arrows, circles, or highlights, to preserve a clean, unbiased evaluation of visual reasoning.
Finally, we construct key rationales that explain the winning variant’s superior performance observed in A/B test results, curated by a group of three experienced UI/UX experts aware of the outcomes. Such expert involvement is essential, as interpreting user behavior often requires deep domain expertise. To minimize the risk of overlooking key behavioral drivers, each rationale was collaboratively developed and refined through multiple rounds of expert review to ensure salience and comprehensive coverage. The process was guided by textual descriptions from data sources as well as well-established principles, including the Nielsen Norman 10 Usability Heuristics (Nielsen 1994) and 12 carefully selected UX laws. These were chosen for their ability to explain how UI elements influence user behavior and capture key aspects of UX. Each rationale is thus tagged with the corresponding law. Since user behavior is often shaped by multiple factors, a single instance may include multiple rationales. See Supplementary for more curation details, including the complete list of UX laws used.
Figure 3: Benchmark construction pipeline. (a) Raw A/B test data were collected from reliable platforms, (b) then UI images were refined for clarity. (c) Finally, essential rationales explaining the test outcomes were crafted by combining insights from both UI/UX experts and data sources.
3.2 Benchmark Analysis
Our pipeline produces a benchmark of 300 real-world UI image pairs, each with a definitive winner identified through A/B test results and annotated with expert-curated key rationales explaining the core design elements that drove user behavior, totalling 684 entries.
Ensuring diversity in our dataset is critical, as effective UI/UX strategies for guiding user behavior vary widely across contexts (Oinas-Kukkonen and Harjumaa 2018). For example, each page type represents a distinct stage in the user journey, often requiring different design strategies. To capture this, WiserUI-Bench spans 11 page types, 11 industry domains, and 2 device types. Common page types include product/service detail pages (29.0%) and landing pages (18.3%). In terms of industry, retail & e-commerce (43.0%) dominates, reflecting its heavy use of A/B testing for conversion optimization, followed by IT & software (13.7%) and travel & hospitality (13.3%). 84% of the UIs are from web platforms and 16% from mobile. Moreover, since our benchmark is based on real-world UIs, it exhibits a wide variety of design changes. We categorize these into 10 high-level types, with a single instance often involve multiple UI changes. In total, we record 490 distinct UI differences across all samples. For example, the UIs shown in Figure 2 reflect changes in layout & structure (repositioning favorite button) and controls & copy (enlarging the ’add to cart’ button). See Supplementary for detailed statistics.
Figure 4: Rationale statistics for WiserUI-Bench: (a) by UX dimensions and laws; (b) by count per instance.
WiserUI-Bench also exhibits wide variation in the types of reasoning used to explain user behavior, as illustrated in Figure 4. Each instance contains an average of 2.28 rationales, with up to 5 in some cases. The UX laws tagged during curation are grouped into three cognitive dimensions of UX—perception, memory, and action—with four laws assigned to each, drawing on the idea from Norman (2007) that user experience can be interpreted through these dimensions. Each rationale is categorized accordingly, enabling fine-grained evaluation of how well models capture the diverse UX factors underlying each design choice. Refer to Supplementary for more details.
4 Task Description
Proper evaluation of a model’s multimodal understanding of UI/UX design requires testing not only whether the model can identify the more effective design in guiding user behavior, but also whether it can explain why, grounding its reasoning in the visual and contextual cues. To this end, we define two independent tasks, decision and explanation, as illustrated in Figure 2, both essential for demonstrating genuine understanding. Success requires identifying key visual differences between two UIs, as well as inferring the likely behavioral goal from context (e.g., encouraging sign-ups) and how the interface supports that goal.
Task 1: UI/UX Design Selection
The first task assesses whether the model can identify which of two UI/UX designs is more effective at influencing user behavior, with the correct answer determined by large-scale A/B testing run by real companies. Although the task is framed as a binary choice, solving it effectively requires nuanced visual reasoning beyond surface-level comparison.
Task 2: UI/UX Design Rationale Alignment
The second task evaluates a model’s reasoning ability in depth. Unlike the first task, it presents the winning UI/UX design upfront and asks the model to explain the factors contributing to its effectiveness in guiding user behavior. This setup mirrors the condition under which expert rationales were made, providing a fair basis for comparison. The model’s explanation is then compared to expert-curated key rationales, and the degree of alignment indicates the quality of its UI/UX reasoning at an expert-level.
5 Experiments
Models
We evaluate a diverse set of MLLMs, spanning both proprietary and open-source models. On the proprietary side, we include o1 (OpenAI et al. 2024b), GPT-4o (OpenAI et al. 2024a), Claude 3.5 Sonnet (Anthropic 2024), all noted for their strong visual reasoning. For open-source models, we evaluate widely adopted models such as Qwen2.5-VL (7B, 32B) (Bai et al. 2025), InternVL-2.5 (8B, 38B) (Chen et al. 2025), LLaVA-NeXT-7B (Liu et al. 2024a) and LLaVA-OneVision-7B (Li et al. 2024b).
5.1 UI/UX Design Selection
Metrics
Model accuracies are reported across input-order variations of each UI image pair, reflecting the multi-image input setting. Specifically, we report: (1) First Accuracy (FA), when the more effective UI/UX design is presented first in the input; (2) Second Accuracy (SA), when it is presented second; and (3) Average Accuracy (AA), the mean of FA and SA, reflecting overall correctness. Finally, we compute (4) Consistent Accuracy (CA), which checks whether the model selects the same, correct UI across both input orders of each pair. This metric isolates the model’s core visual reasoning ability by ensuring that its choices are based on content rather than position bias, a well-known issue in pairwise comparison settings (Wang et al. 2025b; Zhang et al. 2023). (Random guessing yields a CA of 25%: since there’s a 50% chance of being correct in each order.) All evaluations are conducted over three independent runs, and we report the averaged results.
Model FA SA AA CA
Random 50.00 50.00 50.00 25.00
o1 16.56 97.78 57.17 15.56
GPT-4o 31.89 88.33 60.11 30.11
Claude 3.5 Sonnet 26.11 86.67 56.39 24.22
Qwen2.5-VL-32B 33.67 81.44 57.56 31.00
Qwen2.5-VL-7B 23.00 77.89 50.44 12.89
InternVL-2.5-38B 55.67 60.00 57.83 34.56
InternVL-2.5-8B 58.33 45.67 52.00 21.33
LLaVA-NeXT-7B 54.44 40.56 47.50 10.78
LLaVA-OneVision-7B 19.78 81.56 50.67 10.44
Table 2: Experimental results for the UI/UX design selection task on WiserUI-Bench. Best values are in bold; second-best are underlined. All metrics are reported as percentages (%).
Results
Table 2 summarizes model performance on the UI/UX design selection task. All AA scores barely clear the 50% random baseline for a binary choice. Moreover, they are inflated by strong position bias, especially in proprietary models and the Qwen2.5-VL family, which tend to pick the second image. When we examine CA, the order-invariant metric, performance drops back to near-random levels. Interestingly, Qwen2.5-VL-32B and InternVL-2.5-38B score slightly higher than proprietary models on CA, but still fall short of being adequate. Contrary to expectations, o1 shows the strongest bias and a lower CA than GPT-4o, suggesting that self-reasoning training of o1 does not transfer to this task. Smaller (7B/8B) models perform significantly worse than larger ones, confirming the scaling trend. Overall, current MLLMs struggle to consistently select the more effective UI/UX design in real-world UI pairs.
Figure 5: Qualitative results produced by o1, GPT-4o, and Qwen2.5-VL-32B on the two core tasks. (Left) Qwen2.5-VL-32B and GPT-4o consistently selected the correct and incorrect variants, respectively, whereas o1 showed position bias. (Right) o1 successfully captured all key rationales behind the winning design. Green/blue highlights mark alignment with expert rationales.
5.2 UI/UX Design Rationale Alignment
Metrics
To judge how well a model’s generated rationale explains the observed outcome, we compare it to a set of expert-curated key rationales per instance. We report two metrics: Rationale-wise accuracy, which checks model agreement with each expert rationale separately, and Instance-wise accuracy, which counts an instance as correct if the model matches any rationale. Alignment is assessed using a GPT-4o-based evaluator (Liu et al. 2023), validated against human judgments on 1,000 random samples (70.8% accuracy, Cohen’s kappa 0.42). We also report BLEU, ROUGE, METEOR, and BERTScore (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2020) in Supplementary.
Model Rationale-wise Instance-wise
o1 64.18 78.33
GPT-4o 50.15 66.67
Claude 3.5 Sonnet 63.16 78.67
Qwen2.5-VL-32B 60.23 73.00
Qwen2.5-VL-7B 43.71 61.67
InternVL-2.5-38B 45.32 59.33
InternVL-2.5-8B 35.09 51.67
LLaVA-NeXT-7B 11.99 21.67
LLaVA-OneVision-7B 21.78 33.33
Table 3: Results for the UI/UX design rationale alignment task on WiserUI-Bench, figures in percentages (%).
Results
Table 3 reports performance on the UI/UX design rationale alignment task. o1 and Claude 3.5 Sonnet stand out with expert-level explanations, showing strong post-hoc reasoning once the correct answer is given, even though they noticeably fell short on selection when the answer was not avaliable in Task 1. GPT-4o performs worse in this setting, while Qwen2.5-VL-32B remains solid and InternVL-2.5-38B drops. Qwen2.5-VL-7B shows relatively strong alignment, while other 7B/8B models continue to underperform. The consistent gap between instance-wise and rationale-wise scores indicates that while models can often align with part of the expert reasoning, they rarely capture it fully.
6 Analyses
6.1 Qualitative Analysis
We analyzed differences in model outputs in depth. The left part of Figure 5 illustrates an example from the UI/UX design selection task. Variant 1, the actual A/B test winner with larger product images that are more likely to capture users’ attention, was correctly chosen by Qwen2.5-VL-32B, while other models failed to do so. When the correct variant appeared first in the input, Qwen2.5-VL-32B provided appropriate reasoning, demonstrating a solid understanding of UI/UX. In contrast, both o1 and GPT-4o incorrectly attributed their choice to an increased number of items as the reason. Additionally, o1 exhibited position bias.
The right part of Figure 5 shows a case from the UI/UX design rationale alignment task. Given that the UI on the right was the A/B test-verified winner, experts identified the key reasons for its effectiveness as simplified background and enlargened demo button. both of which models must explicitly capture to provide an accurate explanation. o1 succeeds by directly highlighting these key points. GPT-4o correctly identifies the relevant UI elements, but provides an incorrect or vague explanation for their impact. Qwen2.5-VL-32B correctly notes the minimalist design, but stops at identifying the button, missing the key detail of its enlargement and resulting in an incomplete explanation.
6.2 Effect of Diverse Reasoning Strategies
For the first task, we additionally evaluate various reasoning strategies on GPT-4o and Claude 3.5 Sonnet to assess their effectiveness, including Contrastive Chain-of-Thought (CoCoT) (Zhang et al. 2024), Self-Refine (Madaan et al. 2023), Duty-Distinct Chain-of-Thought (DDCoT) (Zheng et al. 2023), and Multi-Agent Debate (MAD) (Liang et al. 2024; Du et al. 2023). For MAD, we stop and moderate the debate after the first (R1) and third (R3) rounds. As shown in Table 4, AA scores remain similar to or below the zero-shot setting, while CA scores often surpass it, especially for MAD (R1) across models, DDCoT on GPT-4o, and CoCoT on Claude 3.5 Sonnet. This gain is mainly due to reduced position bias, enabling more consistent decisions regardless of input order. MAD (R1) performs well by generating and moderating conflicting perspectives, which suits this human preference-driven task, benefiting from multi-perspective reasoning. However, excessive internal debate may hurt performance, as seen in the drop with MAD (R3).
Model Method FA SA AA CA
GPT-4o Zero-Shot 31.89 88.33 60.11 30.11
CoCoT 33.00 83.33 58.17 29.67
Self-Refine 32.44 82.67 57.56 29.89
DDCoT 36.56 86.89 61.72 34.78
MAD (R1) 51.22 67.44 59.33 39.00
MAD (R3) 34.44 80.89 57.67 29.89
Claude 3.5 Sonnet Zero-Shot 26.11 86.67 56.39 24.22
CoCoT 29.78 81.78 55.78 27.22
Self-Refine 25.67 76.44 51.06 20.44
DDCoT 26.22 86.67 56.44 23.56
MAD (R1) 39.00 72.89 55.94 30.00
MAD (R3) 23.33 85.44 54.39 21.44
Table 4: Results for the UI/UX design selection task across reasoning strategies on WiserUI-Bench; metrics match Section 5.1 and are averaged over three runs.
6.3 Rationale Alignment Across UX Aspects
We evaluated rationale-wise accuracy in the second task by each UX dimension, as truly understanding UI/UX requires capturing the full spectrum of cognitive processing—perception, memory, and action—reflected in our rationale categorization. As shown in Figure 6, o1 leads on Perception, indicating strong ability to reason about what users notice in a UI. Claude 3.5 Sonnet performs best on Action, suggesting it is more adept at predicting how users would interact with interfaces. Notably, model performance converge more on Memory than on the others. The persistent sub-70% ceiling across all dimensions underscores the need for deeper cognitive understanding of UX as a whole.
Figure 6: Results for the UI/UX design rationale alignment task by each UX dimension on WiserUI-Bench.
7 Discussion and Conclusion
Toward Behavior-Aware UI Generation: Human Study on Synthetic Pairs
As an exploratory extension of our main study, we evaluated model predictions against a human study on synthetic UI pairs crafted to elicit specific user actions. We position this task and its results as a stepping stone toward automating behavior-aware UI generation.
Prior work on UI generation with large models largely targets matching natural-language or screenshot specifications (Laurençon, Tronchon, and Sanh 2024; Si et al. 2025; Li, Zhang, and Yang 2025) or ensuring basic functionality and visual polish (Wu et al. 2024b; Lu et al. 2025c), rarely examining whether the produced designs actually steer user behavior. Because influencing actions is the core purpose of UI design, generation must also reflect that goal.
Preference-based tuning methods such as Direct Preference Optimization (DPO) (Rafailov et al. 2023) may offer a plausible route. They enable a closed-loop pipeline without direct human supervision: (1) prompt with a behavioral goal, (2) generate candidates, (3) predict which better supports the goal, and (4) feed that signal back to refine the model (Yuan et al. 2024). For this to work, models must first reliably pick the more behavior-effective design not only on real-world interfaces, as in our main study, but also on synthetic pairs.
To test this prerequisite, we conducted a human study with 55 participants on GPT-4o-generated synthetic UI pairs (code and its screenshot) based on 10 behavioral objectives. Participants were shown the objective and chose the variant that better supports it, yielding a preference signal focused on behavior; the majority choice serves as its proxy. Models, in contrast, evaluated the same pairs without access to the goal, mirroring setup in our main study. See Supplementary for details.
Table 5 shows that both proprietary and open-source MLLMs, while most performing better on CA than in main study, still remain near random on AA in this synthetic evaluation, mirroring their limitations observed on WiserUI-Bench. Figure 7 provides a representative example. Going forward, consistently identifying the more effective design in such synthetic pairs will be crucial for behavior-aware UI generation, as it supplies the scalable preference signal needed for models’ closed-loop improvement. We encourage further work in this direction.
Figure 7: Example from the human study. Given the instruction with a behavioral goal, participants selected the better synthetic variant. InternVL2.5-38B failed to match the majority choice due to flawed judgment.
Model FA SA AA CA
GPT-4o 46.67 70.00 58.33 46.67
Claude 3.5 Sonnet 23.33 93.33 58.33 23.33
Qwen2.5-VL-32B 43.33 66.67 55.00 40.00
InternVL-2.5-38B 56.67 46.67 51.67 40.00
Table 5: Results of the models on identifying the behaviorally preferred synthetic UI variant in human study. Reported figures are averaged over three trials.
Conclusion
We introduce a new task that evaluates multimodal understanding of how design drives user actions in the critical field of UI/UX. To support this, we present WiserUI-Bench, a behavior-focused benchmark uniquely built from diverse, real-world A/B-tested UI variants and expert-curated key rationales, providing a reliable basis for assessment. Our two tasks, selection and explanation, go beyond surface-level recognition, requiring models to reason about the behavioral implications of visual design, a cognitively challenging problem. Extensive experiments across both current proprietary and open-source MLLMs reveal they exhibit only shallow levels of nuanced visual reasoning. Achieving stronger performance on this significant task may, for example, require further training on vast amounts of behaviorally grounded UI/UX data that reflect real-world dynamics. We envision WiserUI-Bench and our work as a step toward multimodal systems capable of not only understanding but also generating behavior-aware interfaces, aligning with human cognition and preferences to optimize user experience. More broadly, we hope to catalyze future research toward deeper UI/UX and human-behavior literacy in AI.
A Benchmark for Multimodal UI/UX Design Understanding
Supplementary Material
Appendix A Overview
This supplementary material includes the following sections: (B) Limitations of our work, (C) Ethics Statements for UI/UX experts and human study participants, Detailed descriptions of (D) WiserUI-Bench, (E) experiments, and (F) human study, and concludes with (G) icon attribution.
Appendix B Limitation
Given that user experience patterns vary across cultural contexts and societal norms, cultural biases inevitably exist in our benchmark (Karreman, Romeo, and Li 2016). However, accounting for such diversity is also part of the benchmark’s intent, making it a more realistic and challenging evaluation setting. Also, interactive UIs were not extensively explored in our work. Although the number of 300 UI image pairs may seem limited, publicly available real-world A/B test results are extremely rare, as companies rarely disclose them. Despite the dataset size, it covers a wide range of diverse UI changes, each accompanied by multiple expert-validated rationales.
Appendix C Ethics Statement
To ensure expert-level annotation quality, we recruited three annotators with professional UI/UX backgrounds, all currently employed at global IT companies specializing UX and AI. They were clearly informed of the academic nature of the task and voluntarily participated in the study. Each annotator received monetary compensation equivalent to approximately \$225 USD per person, which reflects fair market payment for domain-specific labor. Payments were issued within 48 hours of annotation completion. All annotators were instructed to focus solely on task-relevant elements and were not exposed to or asked to label any sensitive, offensive, or triggering content.
The human study was also conducted with strict ethical considerations. To minimize biases, we employed a double-blind setup. All 55 participants voluntarily took part in the survey, fully informed that the study was for academic purposes. Additionally, the survey contained no explicit or problematic content, ensuring a safe and respectful research environment.
Appendix D WiserUI-Bench Details
D.1 Image Preprocess
During benchmark construction process, we observed that some UI images from our data sources included manual indicators, such as circled regions, numbered markers, and arrows, highlighting specific UI elements related to reasoning points from A/B test results (see Figure J). While helpful for human interpretation, these indicators introduced artificial attention cues that could potentially bias model evaluation.
To mitigate this, we cleaned the dataset using inpainting techniques. We first identified common indicator patterns via edge detection and color-based segmentation, including circular highlights drawn in semi-transparent colors, numbered markers, and arrows. We then applied automated inpainting followed by manual refinement to remove these elements while preserving UI fidelity. This process ensured a cleaner, more naturalistic dataset for fair visual reasoning evaluation.
D.2 Rationale Annotation
As shown in Figure L, three UI/UX experts are presented with a pair of UI images, one of which was designated as the winner based on A/B test results, along with a corresponding source website link. Experts were instructed to write key rationales explaining why the winning version supported user actions more effectively than the alternative. We engaged experts for annotation, as user behavior is often subconscious and shaped by implicit context, making evaluation by non-experts inherently unreliable. These rationales underwent multiple rounds of review to ensure that no important reason was omitted and no irrelevant reason was included.
Expert annotators were guided by the original sources’ textual description which often highlights key design intentions, and two types of theoretical heuristics selected for their robustness and relevance to user behavior analysis: (1) the Nielsen Norman 10 Usability Heuristics (Nielsen 1994), serving as broad conceptual criteria, and (2) 12 curated UX laws that are still widely applied in practice. These UX laws provide concrete, actionable guidance across diverse aspects of user experience, which can be categorized into three cognitive dimensions: Perception, Memory, and Action. This classification is inspired by the framework proposed by Norman (2007), which introduced the concept of Emotional Design. Emotional design focuses on creating products that evoke emotions and foster positive user experience by engaging users at three cognitive levels: visceral, behavioral, and reflective. The full list of UX laws and their corresponding classifications is presented in Table A.
D.3 Statistics
Figure K presents the distribution of WiserUI-Bench across page types, industry domains, device types, and types of UI element changes. For clarification on the page type classification, the homepage refers to the main entry point of a website, typically offering broad navigation and general content. In contrast, a landing page is usually a standalone page optimized for a specific user action, such as signing up or making a purchase. For clarification on the types of UI element changes, detailed definitions are provided in Table B.
Appendix E Experiment Details
E.1 Setup
Model Configuration
For proprietary MLLMs, we accessed the models via their official API endpoints as follows:
  • o1 (OpenAI et al. 2024b):
    o1-2024-12-17
  • GPT-4o (OpenAI et al. 2024a):
    gpt-4o-2024-08-06
  • Claude 3.5 Sonnet (Anthropic 2024):
    claude-3-5-sonnet-20240620
For open-source models, we used publicly available weights from the Huggingface Hub [^3] https://huggingface.co/models. Specific model versions are as follows:
  • Qwen2.5-VL-32B (Bai et al. 2025):
    Qwen/Qwen2.5-VL-32B-Instruct
  • Qwen2.5-VL-7B (Bai et al. 2025):
    Qwen/Qwen2.5-VL-7B-Instruct
  • InternVL-2.5-38B (Chen et al. 2025):
    OpenGVLab/InternVL2_5-38B
  • InternVL-2.5-8B (Chen et al. 2025):
    OpenGVLab/InternVL2_5-8B
  • LLaVA-NeXT-7B (Liu et al. 2024a):
    llava-hf/llava-v1.6-mistral-7b-hf
  • LLaVA-OneVision-7B (Li et al. 2024b):
    llava-hf/llava-onevision-qwen2-7b-ov-hf
The temperature for all models was set to 0.2.
Resources
All experiments using open-source models were conducted on a Linux machine (Ubuntu 20.04.6 LTS) equipped with 4 NVIDIA RTX A6000 GPUs, each with 48GB VRAM. We used the following software libraries and frameworks: transformers version 4.53.2 and vllm version 0.9.2.
E.2 UI/UX Design Selection
Prompts
Figures A through G show the prompts used, including those designed for diverse reasoning strategies.
Token Counts
Token counts across reasoning strategies on proprietary models are presented in Table C.
Standard Deviation
Table D presents the standard deviation across three experiment runs for each model and reasoning method. For metric abbreviation, FS and SS denote the standard deviation when the first or second UI image input was the more effective one, while AS and CS refer to the standard deviation of total correct and both correct cases, respectively.
Additional Qualitative Results
Figure N and O present the additional qualitative examples across different models. Figure P shows an example comparing various reasoning strategies executed with GPT-4o.
E.3 UI/UX Design Rationale Alignment
Prompts
Figure H presents the prompt used for models’ rationale generation, and Figure I presents the prompt used for its GPT-4o-based evaluation.
Results on Text Similarity Metrics
Alongside rationale-wise and instance-wise accuracies from GPT-4o-based evaluation, we report BLEU, ROUGE, METEOR, and BERTScore as well. (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2020) in Table E.
Additional Qualitative Results
Figure Q presents an additional qualitative example across different models.
Appendix F Human Study Details
Figure M displays the interface used for human evaluation, where participants select which synthetic UI better adheres to the given instructions, including the specified behavioral goal. Table F presents the results of the human study, showing the distribution of participant responses across 10 A/B questions. For each question, the option chosen by the majority is treated as a proxy for human preference and is highlighted in bold. Results of standard deviation across three experiment runs are shown in Table G. Figure R provides an additional qualitative example of how models performed in identifying the preferred UI variant.
Appendix G Icon Attribution
Icons used in this paper are from Flaticon (https://www.flaticon.com), and they are attributed to the respective authors as required by Flaticon’s license.
UX Dimension UX Law Description
Perception Law of Common Region (Koffka 1922) Elements sharing a boundary or background are perceived as a unified group. This can be a visible border, a colored box, or a distinct region in the layout.
Law of Proximity (Koffka 1922) Items positioned close together are seen as belonging to the same group. This mental shortcut helps users quickly identify relationships between adjacent elements.
Law of Prägnanz (Koffka 1922) Humans naturally perceive and interpret complex images in their simplest forms. It reduces cognitive effort by helping us group and understand elements more quickly.
Law of Similarity (Koffka 1922) People group together elements that look alike, assuming they serve a similar purpose. Consistency in color, shape, or size leads the user to see these items as related.
Memory Von Restorff Effect (Von Restorff 1933) People remember a distinct element more easily than others that blend in. This is also known as the Isolation Effect, emphasizing our attention on anything unusual.
Serial Position Effect (Ebbinghaus, Ruger, and Bussenius 1913) Users tend to remember the first and last items in a series more than those in the middle. This influences how they recall sequential information or actions.
Miller’s Law (Miller 1956) On average, people can hold only about seven items in their working memory at once. Going beyond this limit causes cognitive overload.
Jakob’s Law (Nielsen 2000) Users expect new interfaces to operate similarly to those they’ve used before. Leveraging established patterns lowers cognitive load and accelerates adoption.
Action Hick’s Law (Hick 1952) Decision time rises with the number and complexity of choices. Too many options can overwhelm users and slow their actions.
Pareto Principle (Pareto 1919) Roughly 80% of outcomes stem from 20% of causes. Focusing on the most impactful features or tasks delivers the greatest improvement.
Fitts’s Law (Fitts and Peterson 1964) The time required to reach a target depends on its size and distance. Larger, closer elements are easier to access quickly and accurately.
Goal-Gradient Effect (Hull 1932) Motivation increases as people get closer to finishing a task. Providing clear progress markers encourages users to continue and complete it faster.
Table A: Overview of UX laws used for rationale annotation, and their categorization by cognitive dimension (Perception, Memory, Action).
Type Definition
Content & Copy All textual elements that convey meaning or guide the user, including headlines, body text, labels, error messages, promotional language, and social proof content.
Visual & Style Visual design elements such as color schemes, typography, spacing, icon sizes, and overall aesthetic consistency that shape the look and feel of the interface.
Controls & Inputs Interactive UI components that allow users to perform actions, including buttons, toggles, dropdowns, sliders, search fields, and other input elements.
Layout & Structure The spatial organization of UI components, covering alignment, grouping, column or grid usage, visual hierarchy, and structural consistency across screens.
Navigation & Flow Interface elements that support user movement and orientation within the product, such as menus, sidebars, breadcrumbs, tabs, and filter panels.
Media & Visual Assets Non-textual content like images, icons, thumbnails, banners, and videos used to support communication, branding, or user engagement.
Pricing & Promotion How pricing details, discounts, promotional messages, limited-time offers, or benefits like free trials are visually and textually communicated to the user.
Trust & Social Proof UI elements that aim to build user trust and credibility, including customer reviews, ratings, trust badges, security seals, guarantees, and testimonials.
Forms & Form UX Design and usability aspects of data entry forms, such as number and arrangement of fields, required indicators, input validation, and overall ease of completion.
Status & Feedback Visual and textual responses to user actions or system states, including loading indicators, progress bars, tooltips, error messages, confirmations, and alerts.
Table B: Definitions of UI element change types.
Models Methods Input Tokens Output Tokens
o1 Zero-Shot 1292.5 ($\pm$ 559.77) 1347.7 ($\pm$ 489.76)
GPT-4o Zero-Shot 1453.4 ($\pm$ 634.41) 248.8 ($\pm$ 57.98)
CoCoT 1458.4 ($\pm$ 634.41) 329.7 ($\pm$ 63.29)
Self-Refine 5240.4 ($\pm$ 1935.80) 820.9 ($\pm$ 142.43)
DDCoT 1521.4 ($\pm$ 634.41) 415.3 ($\pm$ 73.42)
MAD (R1) 4942.4 ($\pm$ 1903.15) 483.1 ($\pm$ 58.5)
MAD (R3) 11501.5 ($\pm$ 4435.48) 1158.1 ($\pm$ 82.60)
Claude 3.5 Sonnet Zero-Shot 1686.6 ($\pm$ 1062.43) 401.4 ($\pm$ 72.94)
CoCoT 1692.5 ($\pm$ 1062.58) 487.0 ($\pm$ 82.28)
Self-Refine 6230.7 ($\pm$ 3245.69) 1201.4 ($\pm$ 207.32)
DDCoT 1754.5 ($\pm$ 1064.29) 582.2 ($\pm$ 97.13)
MAD (R1) 5895.5 ($\pm$ 3224.31) 834.1 ($\pm$ 125.13)
MAD (R3) 13853.4 ($\pm$ 7512.63) 1936.5 ($\pm$ 276.14)
Table C: Overview of Token Counts ($\mu\pm\sigma$) in the UI/UX design selection task on WiserUI-Bench.
Model Method FS SS AS CS
o1 Zero-Shot 1.528 2.082 3.606 0.577
GPT-4o Zero-Shot 1.528 1.000 2.082 1.528
CoCoT 4.000 2.646 1.732 1.732
Self-Refine 2.517 1.000 3.215 0.577
DDCoT 2.082 2.309 0.577 2.517
MAD (R1) 3.215 3.055 6.245 2.646
MAD (R3) 0.577 2.082 1.732 3.215
Claude 3.5 Sonnet Zero-Shot 1.528 2.000 3.512 2.887
CoCoT 3.055 3.512 0.577 0.577
Self-Refine 2.646 3.512 4.619 1.528
DDCoT 2.082 1.732 0.577 2.517
MAD (R1) 1.000 1.528 2.309 4.359
MAD (R3) 1.000 2.887 3.786 3.786
Qwen2.5-VL-32B Zero-Shot 1.732 2.309 2.082 2.000
Qwen2.5-VL-7B Zero-Shot 2.646 0.577 2.517 3.215
InternVL-2.5-38B Zero-Shot 2.646 1.000 3.000 3.055
InternVL-2.5-8B Zero-Shot 1.000 2.646 2.646 2.000
LLaVA-NeXT-7B Zero-Shot 3.512 1.155 4.583 6.028
LLaVA-OneVision-7B Zero-Shot 4.041 2.309 4.359 3.512
Table D: Results on standard deviation across three experiment runs in the UI/UX design selection task for each model and reasoning strategy on WiserUI-Bench.
Model BLEU-2 BLEU-4 ROUGE-L METEOR BERTScore
o1 0.0181 0.0041 0.0972 0.1630 0.8414
GPT-4o 0.0162 0.0040 0.0875 0.1557 0.8233
Claude 3.5 Sonnet 0.0141 0.0037 0.0654 0.1418 0.8076
Qwen2.5-VL-32B 0.0098 0.0025 0.0442 0.1016 0.8085
Qwen2.5-VL-7B 0.0180 0.0048 0.0789 0.1481 0.8331
InternVL-2.5-38B 0.0128 0.0034 0.0574 0.1225 0.8080
InternVL-2.5-8B 0.0111 0.0032 0.0510 0.1124 0.8082
LLaVA-NeXT-7B 0.0077 0.0018 0.0386 0.0908 0.8127
LLaVA-OneVision-7B 0.0157 0.0043 0.0718 0.1416 0.8255
Table E: Results of text similarity metrics per expert rationale for the UI/UX design rationale alignment task on WiserUI-Bench.
Choice Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q10
A 42 9 23 14 25 40 14 27 35 18
B 13 46 32 41 30 15 41 28 20 37
Majority Choice A B B B B A B B A B
Table F: Distribution of responses from human participants indicating their preferred UI variant across 10 A/B test questions.
Model FS SS AS CS
GPT-4o 0.577 0.000 0.577 0.577
Claude 3.5 Sonnet 0.577 0.577 0.577 0.577
Qwen2.5-VL-32B 0.577 0.577 1.000 0.000
InternVL-2.5-38B 0.577 0.577 0.577 1.000
Table G: Results on standard deviation across three experiment runs on identifying the behaviorally preferred synthetic UI variant in human study.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. Identify the key UI differences between the two versions, and then evaluate which variant is more effective UI/UX design that leads to better user experience and conversion. You should end your answer with following the format (No bold, etc): More effective: <First/Second>
Figure A: Prompt used for the UI/UX design selection task with zero-shot strategy.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. Identify the key UI differences between the two versions, and then evaluate which variant is more effective UI/UX design that leads to better user experience and conversion. Think step by step. You should end your answer with the following format (No bold, etc): More effective: <First/Second>
Figure B: Prompt used for the UI/UX design selection task with CoCoT strategy.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. We want to identify the key UI differences between the two versions, and then evaluate which variant is more effective UI/UX design that leads to better user experience and conversion. Review your previous answer and find problems with your answer. Previous answer: {previous_answer}
Figure C: Prompt used for the UI/UX design selection task with Self-Refine strategy on the reviewing step. Prompt inputs are in bold.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. We want to identify the key UI differences between the two versions, and then evaluate which variant is more effective UI/UX design that leads to better user experience and conversion. Your previous answer: {previous_answer} Feedback on your previous answer: {feedback} Based on the feedback for your previous answer, improve your answer. You should end your answer with following the format (No bold, etc): More effective: <First/Second>
Figure D: Prompt used for the UI/UX design selection task with Self-Refine strategy on the improving step. Prompt inputs are in bold.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. We want to identify the key UI differences between the two versions and then evaluate which variant is more effective UI/UX design that leads to better user experience and conversion. You should (1) Think step by step about the preliminary knowledge to answer the question, deconstruct the problem as completely as possible down to necessary sub-questions. (2) With the aim of helping answer the original question, try to answer the sub-questions. (3) Give your final answer according to the sub-questions and sub-answers. You should end your answer with the following format (No bold, etc): More effective: <First/Second>
Figure E: Prompt used for the UI/UX design selection task with DDCoT strategy.
You are an expert in designing UI/UX for web/apps. And you are a debater. Hello and welcome to the debate. The two screenshots show two different versions of the same page. Identifying the key UI differences between the two versions, you think the {first/second} variant is more effective UI/UX design that leads to better user experience and conversion. It’s not necessary to fully agree with each other’s perspectives, as our objective is to find the correct answer. The opponent proposed that: {opponent_opinion}
Figure F: Prompt used for the UI/UX design selection task with MAD strategy on the debate step arguing the first/second version better drives user actions. Prompt inputs are in bold.
You are an expert in designing UI/UX for web/apps. And you are a moderator. There will be two debaters involved in a debate. The two screenshots show two different versions of the same page. They are discussing which variant is more effective UI/UX design that leads to better user experience and conversion. At the end of each debate round, you will evaluate answers and decide which is correct. Now the {number} round of debate for both sides has ended. Argue from the side arguing for the first version: {first_reason} Argue from the side arguing for the second version: {second_reason} Then what is the correct answer? Please give the final answer that you think is correct and summarize your reasons. You should end your answer with following the format (No bold, etc): More effective: <First/Second>
Figure G: Prompt used for the UI/UX design selection task with MAD strategy on the moderation step. Prompt inputs are in bold.
You are an expert in designing UI/UX for web/apps. The two screenshots show two different versions of the same page. If the SECOND version has proven to be more effective UI/UX design that leads to better user experience and conversion, provide all possible and appropriate for its superiority.
Figure H: Prompt used for the UI/UX design rationale alignment task.
Given two statements are reasons why one UI is better than another in terms of guiding user behavior effectively. Statement 1: {model_reason}
Statement 2: {reference_reason} Your task is to assess whether the first statement contains the same reasoning and essential content as the second statement. If it does, answer with ’Yes’. Otherwise, answer with ’No’.
Figure I: Prompt used for GPT-4o-based evaluation on the UI/UX design rationale alignment task. Prompt inputs are in bold.
Figure J: Example of visual indicator removal in WiserUI-Bench. Left: Original image with markers added by data source website. Right: After inpainting, with markers removed while preserving UI elements.
Figure K: Detailed statistics of WiserUI-Bench on (a) page type, (b) industry domain, (c) device type, (d) UI element change type
Figure L: A screenshot of the rationale annotation interface for UI/UX experts based on original data sources and guidelines. While writing the key rationales, they also selected the corresponding UX law.
Figure M: A screenshot of the human study interface. Each participant selects the UI variant they believe better supports the given instruction, including the specified behavioral goal.
Figure N: A qualitative example from the UI/UX design selection task on WiserUI-Bench: Both Claude 3.5 Sonnet and InternVL-2.5-38B selected the A/B test winner.
Figure O: A qualitative example from the UI/UX design selection task on WiserUI-Bench: GPT-4o selected the A/B test winner, whereas LLaVA-NeXT-7B did not.
Figure P: A qualitative example from the UI/UX design selection task on WiserUI-Bench, showcasing the effect of diverse reasoning strategies executed on GPT-4o.
Figure Q: A qualitative example from the UI/UX design rationale alignment task on WiserUI-Bench: InternVL-2.5-8B captured only the memory rationale, while Qwen2.5-VL-32B identified two key rationales. The highlights in the models’ responses indicate the parts that align with the expert-curated key rationales.
Figure R: A qualitative example from the UI/UX design selection on human study: InternVL2.5-38B failed to match the majority choice, whereas Claude 3.5 Sonnet succeeded. For fair evaluation, we inserted the same images into the placeholder area of each generated UI, as shown above.

### v4

@@
 Abstract
-User interface (UI) design goes beyond visuals, guiding user behavior and overall user experience (UX).
-Strategically crafted interfaces, for example, can boost sign-ups and drive business sales, underscoring the shift toward UI/UX as a unified design concept.
-While recent studies have explored UI quality evaluation using Multimodal Large Language Models (MLLMs), they largely focus on surface-level features, overlooking behavior-oriented aspects.
-To fill this gap, we introduce WiserUI-Bench  [^1] https://github.com/jeochris/wiserui-bench, a novel benchmark for assessing models’ multimodal understanding of UI/UX design.
-It includes 300 diverse real-world UI image pairs, each consisting of two design variants A/B-tested at scale by actual companies, where one was empirically validated to steer more user actions than the other.
-Each pair is accompanied one or more of 684 expert-curated rationales that capture key factors behind each winning design’s effectiveness, spanning diverse cognitive dimensions of UX.
-Our benchmark supports two core tasks: (1) selecting the more effective UI/UX design by predicting the A/B test verified winner and (2) assessing how well a model, given the winner, can explain its effectiveness in alignment with expert reasoning.
-Experiments across several MLLMs show that current models exhibit limited nuanced reasoning about UI/UX design and its behavioral impact.
-We believe our work will foster research in UI/UX understanding and enable broader applications such as behavior-aware interface optimization.
+User interface (UI) design goes beyond visuals to shape user experience (UX), underscoring the shift toward UI/UX as a unified concept.
+While recent studies have explored UI evaluation using Multimodal Large Language Models (MLLMs), they largely focus on surface-level features, overlooking how design choices influence user behavior at scale.
+To fill this gap, we introduce WiserUI-Bench  [^1] https://github.com/jeochris/wiserui-bench , a novel benchmark for multimodal understanding of how UI/UX design affects user behavior, built on 300 real-world UI image pairs from industry A/B tests, with empirically validated winners that induced more user actions.
+For future design progress in practice, post-hoc understanding of why such winners succeed with mass users is also required; we support this via expert-curated key interpretations for each instance.
+Experiments across multiple MLLMs on WiserUI-Bench for two main tasks, (1) predicting the more effective UI image between an A/B-tested pair, and (2) explaining it post-hoc in alignment with expert interpretations, show that models exhibit limited understanding of the behavioral impact of UI/UX design.
+We believe our work will foster research on leveraging MLLMs for visual design in user behavior contexts.
+Do MLLMs Capture How Interfaces Guide User Behavior?
+A Benchmark for Multimodal UI/UX Design Understanding
+Jaehyun Jeon1, Min Soo Kim1, Jang Han Yoon1, Sumin Shim1, Yejin Choi1, Hanbin Kim1, Dae Hyun Kim1thanks: Co-corresponding authors, Youngjae Yu2footnotemark: 1Yonsei University  2Seoul National University {jaehyun.jeon, dhkim16}@yonsei.ac.kryoungjaeyu@snu.ac.kr
 1 Introduction
 User interface (UI) design is central to application development.
-While visual aesthetics lay its foundation, the fundamental goal is to steer user behavior during interaction, such as encouraging sign-ups or purchases, and thereby boost service revenue (Fogg 2002).
-This behavioral focus expands to UI/UX design, uniting visual form with user experience (UX) considerations (Norman 1988).
-Figure 1: A real-world behavior-driven design decision example from our benchmark, grounded in actual A/B test results.
-Such cases reflect UI changes that steer user actions, making them effective testbeds for evaluating models’ visual reasoning about UI/UX design.
-Practitioners validate such decisions through large-scale A/B tests that randomly assign users to alternative UI variants and measure which one more effectively drives the desired action (Kohavi et al.
-2009).
-These tests often reveal that even subtle design changes can significantly lift conversions.
-For example, as shown in Figure 1, simply reducing the number of options lowered user friction and increased sign-ups.
-Given that predicting and explaining such outcomes is nontrivial even for humans, hinging on complex cognitive processes of user experience and implicit human preferences, assessing multimodal understanding of UI/UX design becomes a crucial challenge, demanding not only identifications of visual elements but also reasoning about how design influences user behavior.
-Benchmark # Samples Unit Source Quality Source Type
-Large-Scale User
-Validation
-Expert
-Rationale
-Evaluation Objective
-Yang and Li (2024)
-382 Single Legacy UI Mobile
-Guideline Violation
-Detection
-JitterWeb
-(Wu et al. 2024a)
-2.3M Pair Synthetic Mobile + Web
-Basic Visual Quality
-BetterWeb
-(Wu et al. 2024a)
-892 Pair Synthetic Mobile
-Basic Visual Quality
-UICrit
-(Duan et al. 2024)
-983 Single Legacy UI Mobile
-UI Design Critique
-WiserUI-Bench
-300 Pair
-Actual
-Production
-Mobile + Web
-(Real A/B Test)
-UI/UX Design Understanding
-(User Behavior-Oriented)
+While visual aesthetics lay its foundation, the fundamental goal is to steer user behavior during interaction, such as encouraging sign-ups or purchases, and thereby boost service revenue (Fogg, 2002).
+This behavioral focus expands to UI/UX design, uniting visual form with user experience (UX) considerations (Norman, 1988).
+Figure 1: A real-world, behavior-aware design decision example from WiserUI-Bench, grounded in A/B test results, illustrating how UI changes steer user actions.
+Practitioners validate such decisions through large-scale A/B tests that randomly assign users to alternative UI variants and measure which design more effectively drives the desired action (Kohavi et al., 2009).
+For example, as shown in Figure 1, version B achieves higher sign-ups overall, interpreted as reduced user friction from fewer options.
+Such interpretation reflects an abductive reasoning process (Hobbs et al., 1993; Bhagavatula et al., 2020), inferring the most plausible explanations for aggregate outcomes that are used to guide subsequent design decisions (Quin et al., 2024).
+Given that predicting and explaining these results is nontrivial even for humans, hinging on complex cognitive processes underlying user experience, assessing multimodal understanding of UI/UX design becomes a crucial challenge, requiring not only identifying visual elements but also reasoning about how they influence user behavior.
+Figure 2: Overview of WiserUI-Bench and two main tasks.
+Each instance contains a UI image pair with verified A/B test winner and expert-curated key interpretations explaining it.
+These span the three cognitive UX dimensions; the example includes all three, though most cover fewer.
+MLLMs are evaluated on (1) selecting the more effective UI/UX design by predicting the verified winner, and (2) explaining the effectiveness of a given winner, measured by whether the model captures each expert interpretation.
+Recent efforts have explored how to evaluate UI design quality with Multimodal Large Language Models (MLLMs); however, they all fall short in capturing the behavioral impact of UI/UX design, as summarized in Table 1.
+While some focus solely on basic visual, heuristic-based attributes (Yang and Li, 2024; Wu et al., 2024), others rely on handwritten critiques of individual UIs, lacking validation against real user data and offering limited coverage of cognitive aspects (Duan et al., 2024).
+To address these limitations, we introduce WiserUI-Bench (Pairwise UI/UX Design Understanding Benchmark), a novel benchmark for evaluating a model’s multimodal understanding of how UI/UX design affects user behavior.
+WiserUI-Bench comprises 300 real-world UI image pairs with winning variants verified by A/B tests from actual companies, spanning varied contexts and sourced from reliable platforms.
+Each pair includes expert-curated key interpretations, 684 in total, written post-hoc with knowledge of the winner to explain its effectiveness, covering core behavioral drivers.
+An example instance is illustrated in the upper part of Figure 2.
+Grounded in large-scale real user behavior data, our benchmark captures authentic usage and preference patterns, providing a reliable basis for evaluation.
+Each interpretation is tagged with one of the three cognitive UX dimensions articulated by Norman (2007): perception (what users notice), memory (what users remember), and action (how users act).
+Our benchmark supports two core tasks, illustrated in the lower part of Figure 2: (1) selecting the more effective UI/UX design from a given pair by predicting the A/B test-verified winner, and (2) interpreting the UI/UX effectiveness of the given winner by assessing alignment between model-generated and expert-curated interpretations.
+This task setup captures the essential capabilities for UI/UX assessment, not only to identify which design better guides user behavior, but also to reliably explain why in a post-hoc manner.
+Experiments on a broad range of proprietary and open-source MLLMs on WiserUI-Bench show that existing models struggle with visual reasoning in user behavior contexts.
+Models perform near-random on selection, and their post-hoc interpretations remain insufficient to reach expert level.
+We expect that WiserUI-Bench and our novel UI/UX-focused tasks will spur further research on this critical and underexplored, user-centered area, advancing behavior-aware reasoning about visual design in MLLMs.
+Benchmark # Samples Unit Source Quality Source Type Large-Scale User Validation Expert Interpretation Evaluation Objective
+Yang and Li (2024) 382 Single Legacy UI Mobile Guideline Violation Detection
+BetterWeb (Wu et al., 2024) 892 Pair Synthetic Mobile Basic Visual Quality
+UICrit (Duan et al., 2024) 983 Single Legacy UI Mobile UI Design Critique
+WiserUI-Bench (Ours) 300 Pair Actual Production Mobile + Web (Real A/B Test) UI/UX Design Understanding (User Behavior-Focused)
 Table 1: Comparison between WiserUI-Bench and existing UI evaluation benchmarks.
-Recent efforts have explored how to evaluate UI design quality with Multimodal Large Language Models (MLLMs); however, they all fall short in capturing the behavioral impact of UI/UX design, as summarized in Table 1.
-While some focus solely on basic visual, heuristic-based attributes (Yang and Li 2024; Wu et al.
-2024a), others rely on handwritten critiques of individual UIs, lacking validation against real user behavior data and offering limited coverage of cognitive aspects (Duan et al.
-2024).
-To address these limitations, we introduce WiserUI-Bench (Pairwise UI/UX Design Understanding Benchmark), a novel benchmark for evaluating a model’s multimodal understanding of UI/UX design.
-WiserUI-Bench includes 300 real-world UI image pairs with winning variants verified by A/B tests from actual companies, spanning varied contexts and sourced from reliable platforms.
-Each pair is accompanied by one or more expert-curated rationales, 684 in total, written post hoc with knowledge of the winner to explain its effectiveness, covering key behavioral drivers.
-An example instance is illustrated in the upper part of Figure 2.
-Uniquely grounded in large-scale real user behavior data, our benchmark mirrors authentic usage and preference patterns, providing a reliable basis for evaluation.
-Each rationale is tagged with one of the three cognitive UX dimensions articulated by Norman (2007)—perception, what users notice; memory, what they remember; and action, how they act—and with the specific UX law it reflects.
-This enables fine-grained assessment of how well models grasp the underlying user experience behind each design choice.
-For example, the action-related rationale in Figure 2 corresponds to Fitts’s Law, which states that users act faster when targets are larger and closer (Fitts and Peterson 1964).
-Our benchmark supports two core tasks, illustrated in the lower part of Figure 2: (1) selecting the more effective UI/UX design from a given pair by predicting the A/B test verified winner, and (2) assessing how well model-generated rationales, given the winner, align with expert reasoning.
-This dual-task setup captures the essential capabilities for UI/UX assessment, not only identifying which design better guides user behavior, but also explaining why reliably.
-Experiments on a broad range of proprietary and open-source MLLMs on WiserUI-Bench show that current models struggle with behavior-centered visual reasoning.
-Models perform near-random on selection, and their post-hoc reasoning remains insufficient to reach expert level.
-We expect that WiserUI-Bench and our novel UI/UX-focused tasks will spur further research on this critical problem and potentially push toward behavior-aware UI generation, the direction that truly matters for achieving user-centered design.
-In summary, our main contributions are as follows:
-  1. 1.
+In summary, our main contributions are:
   • We introduce a novel and challenging task of multimodal UI/UX understanding that centers on the behavioral impact of design.
-  2. 2.
-    We release WiserUI-Bench, a benchmark of 300 real-world UI pairs uniquely constructed from verified results of actual large-scale A/B tests conducted by real companies, along with expert-curated key rationales.
-  3. 3.
-    Through comprehensive experiments and analyses across diverse MLLMs, we uncover their limited ability to reason about behavior-driven UI effectiveness, highlighting a critical gap in current visual reasoning.
-2 Related Works
-Visual Reasoning
+  • We release WiserUI-Bench, a benchmark of 300 real-world UI pairs uniquely constructed from verified results of large-scale industry A/B tests, along with 684 carefully curated expert key interpretations.
+  • Through comprehensive experiments and analyses across diverse MLLMs, we uncover their limited ability to understand the behavioral impact of UI/UX design, highlighting a critical gap in current visual reasoning.
+2 Related Work
+2.1 Visual Reasoning
 Visual reasoning benchmarks for Multimodal Large Language Models (MLLMs) evaluate models’ ability to process visual information and reason at a cognitive level.
-While some target general tasks (Yue et al.
-2024; Liu et al.
-2024b; Fu et al.
-2024; Li et al.
-2024a), others focus on specific domains like mathematics or science (Hao et al.
-2025; Lu et al.
-2024; Xu et al.
-2025).
-However, few benchmarks address reasoning about human preferences in visual contexts.
-Moreover, although some studies explore visual reasoning across multiple images (Cheng et al.
-2025; Zhao et al.
-2024), such cases remain rare.
-Our task incorporates both aspects, reasoning over paired UI images and aligning with behavior-driven preferences, as captured by real-world A/B test outcomes.
-Figure 2: Overview of WiserUI-Bench and two main tasks.
-Each instance contains a UI image pair with verified A/B test winner and expert-curated key rationales explaining it.
-These span the three cognitive UX dimensions; the example includes all three, though most cover fewer.
-MLLMs are evaluated on (1) selecting the more effective UI/UX design by predicting the verified winner, and (2) explaining the effectiveness of a given winner, measured by whether the model covers each expert rationale.
-UI Evaluation
+While some target general tasks (Yue et al., 2024; Liu et al., 2024b; Fu et al., 2024; Li et al., 2024b), others focus on specific domains like mathematics or science (Hao et al., 2025; Lu et al., 2024; Xu et al., 2025).
+However, few benchmarks address reasoning about human behavior or preference patterns in visual contexts.
+Moreover, although some studies explore visual reasoning across multiple images (Cheng et al., 2025; Zhao et al., 2024), such cases remain rare.
+Our task incorporates both aspects, reasoning over paired UI images and aligning with behavior patterns, as captured by actual A/B test outcomes.
+2.2 UI Evaluation
 Recent advances have driven progress in evaluating UI quality with MLLMs.
-Yang and Li (2024) proposes a benchmark for detecting guideline violations in individual screens, while Wu et al.
-(2024a) generates synthetic UI pairs via perturbations such as color noise or spacing changes.
+Yang and Li (2024) propose a benchmark for detecting guideline violations in individual screens, while Wu et al.
+(2024) generates synthetic UI pairs via perturbations such as color noise or spacing changes.
 However, both are limited to heuristic analysis of surface-level appearance.
 Duan et al.
-(2024) evaluates single screens using expert-critique annotations, but the benchmark has key limitations: it lacks validation with real-world user behavior and relies on RICO (Deka et al.
-2017), an outdated, mobile-only dataset that limits its relevance to today’s diverse, production-level designs.
-Prior work also suggests that pairwise evaluation yields more reliable signals, particularly for subjective tasks like ours that target behavior-driven preferences captured by real-world A/B tests (Abel et al.
-2020; Bawabe et al.
-2021).
-In this context, WiserUI-Bench offers a stronger foundation by grounding design effectiveness in large-scale user behavior data and requiring models to engage in deeper, nuanced reasoning over both visuals and user experience in a pairwise format.
-Simulation of Human Behavior
+(2024) evaluates single screens using expert-critique annotations, but it lacks validation with real-world user behavior and relies on RICO (Deka et al., 2017), an outdated, mobile-only dataset that limits its relevance to today’s production-level designs.
+Prior work also suggests that pairwise evaluation yields more reliable signals, particularly for subjective tasks like ours that target behavior patterns captured by real-world A/B tests (Abel et al., 2020; Bawabe et al., 2021).
+In this context, WiserUI-Bench offers a stronger foundation by grounding design effectiveness in large-scale user behavior data and requiring models to engage in deeper, nuanced reasoning over both visuals and UX in a pairwise format.
+2.3 Simulation of Human Behavior
 Efforts to simulate human behavior with AI are ongoing.
@@
 (2025b) and Wang et al.
-(2025a) deploy persona-based agents to browse interfaces and examine how designs may steer user behavior; the former checks realism only via qualitative interviews, while the latter compares against actual A/B tests but in a single scenario.
-By contrast, our work grounds evaluation on numerous A/B test results from real users and introduces a novel multimodal task: reason over UI variant pairs to predict user behavior, enabling direct measurement of behavioral fidelity.
+(2025a) deploy persona-based agents to browse interfaces and examine how designs may steer user behavior; the former checks realism only via qualitative interviews, while the latter compares against an actual A/B test but only under a single scenario.
+By contrast, our work grounds evaluation in numerous real-user A/B test results and introduces diverse multimodal tasks that reason over UI variant image pairs to predict and explain user behavior.
 3 WiserUI-Bench
@@
 By grounding our benchmark in such real, rigorously tested cases, we ensure a credible basis for evaluating UI/UX design understanding in practical scenarios.
+Figure 3: Distribution of (a) UI change element types and (b) expert interpretations for WiserUI-Bench.
+Figure 4: Benchmark construction pipeline.
+(a) Raw A/B test data were collected from reliable platforms, (b) then UI images were refined for clarity.
+(c) Finally, key interpretations explaining the test outcomes were curated by UI/UX experts.
 Data Curation Pipeline
@@
 First, we extract raw UI images and accompanying textual descriptions from our selected platform sources.
-Next, we refine these images by eliminating any visual markers, such as arrows, circles, or highlights, to preserve a clean, unbiased evaluation of visual reasoning.
-Finally, we construct key rationales that explain the winning variant’s superior performance observed in A/B test results, curated by a group of three experienced UI/UX experts aware of the outcomes.
+Next, we refine these images by eliminating any visual markers that were externally added, such as arrows or circles, to preserve a clean, unbiased evaluation of visual reasoning.
+Finally, we construct key interpretations explaining the winning variant’s superior performance observed in A/B test results, curated by three experienced UI/UX experts aware of the outcomes.
 Such expert involvement is essential, as interpreting user behavior often requires deep domain expertise.
-To minimize the risk of overlooking key behavioral drivers, each rationale was collaboratively developed and refined through multiple rounds of expert review to ensure salience and comprehensive coverage.
-The process was guided by textual descriptions from data sources as well as well-established principles, including the Nielsen Norman 10 Usability Heuristics (Nielsen 1994) and 12 carefully selected UX laws.
-These were chosen for their ability to explain how UI elements influence user behavior and capture key aspects of UX.
-Each rationale is thus tagged with the corresponding law.
-Since user behavior is often shaped by multiple factors, a single instance may include multiple rationales.
-See Supplementary for more curation details, including the complete list of UX laws used.
-Figure 3: Benchmark construction pipeline.
-(a) Raw A/B test data were collected from reliable platforms, (b) then UI images were refined for clarity.
-(c) Finally, essential rationales explaining the test outcomes were crafted by combining insights from both UI/UX experts and data sources.
+To identify core behavioral drivers, only interpretations independently identified by at least two experts were retained.
+This resulted in final 684 (562 with full, 122 with two-way agreement) out of 845 initial candidates.
+The process was guided by textual descriptions from data sources and well-established principles, the Nielsen Norman 10 Usability Heuristics (Nielsen, 1994).
+Each was then tagged with the most relevant law from a curated set of 12 UX laws.
+Since user behavior is often shaped by multiple factors, a single instance may include multiple interpretations.
+See Appendix A.1 for details.
 3.2 Benchmark Analysis
-Our pipeline produces a benchmark of 300 real-world UI image pairs, each with a definitive winner identified through A/B test results and annotated with expert-curated key rationales explaining the core design elements that drove user behavior, totalling 684 entries.
-Ensuring diversity in our dataset is critical, as effective UI/UX strategies for guiding user behavior vary widely across contexts (Oinas-Kukkonen and Harjumaa 2018).
-For example, each page type represents a distinct stage in the user journey, often requiring different design strategies.
-To capture this, WiserUI-Bench spans 11 page types, 11 industry domains, and 2 device types.
-Common page types include product/service detail pages (29.0%) and landing pages (18.3%).
-In terms of industry, retail & e-commerce (43.0%) dominates, reflecting its heavy use of A/B testing for conversion optimization, followed by IT & software (13.7%) and travel & hospitality (13.3%).
-84% of the UIs are from web platforms and 16% from mobile.
-Moreover, since our benchmark is based on real-world UIs, it exhibits a wide variety of design changes.
-We categorize these into 10 high-level types, with a single instance often involve multiple UI changes.
-In total, we record 490 distinct UI differences across all samples.
-For example, the UIs shown in Figure 2 reflect changes in layout & structure (repositioning favorite button) and controls & copy (enlarging the ’add to cart’ button).
-See Supplementary for detailed statistics.
-Figure 4: Rationale statistics for WiserUI-Bench: (a) by UX dimensions and laws; (b) by count per instance.
-WiserUI-Bench also exhibits wide variation in the types of reasoning used to explain user behavior, as illustrated in Figure 4.
-Each instance contains an average of 2.28 rationales, with up to 5 in some cases.
-The UX laws tagged during curation are grouped into three cognitive dimensions of UX—perception, memory, and action—with four laws assigned to each, drawing on the idea from Norman (2007) that user experience can be interpreted through these dimensions.
-Each rationale is categorized accordingly, enabling fine-grained evaluation of how well models capture the diverse UX factors underlying each design choice.
-Refer to Supplementary for more details.
+Our pipeline produces a benchmark of 300 real-world UI image pairs, each with a definitive winner identified through A/B test results and annotated with expert-curated key interpretations of the core design factors driving user behavior, totaling 684 entries.
+We include detailed statistics in Appx.
+A.2.
+Ensuring contextual diversity is essential, as effective UI/UX strategies vary widely across use cases (Oinas-Kukkonen and Harjumaa, 2018).
+For example, page types correspond to different stages of the user journey, requiring distinct strategies.
+To capture this, WiserUI-Bench spans 11 page types, 13 industry domains, and two device types.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2: Results of the UI/UX design selection task on WiserUI-Bench, shown overall and by UI change element types.
+Best and second-best model scores for each metric are marked in bold and underlined, respectively.
+For CA, the best and second-best UI element type performances within each model are highlighted in blue and yellow, respectively.
+All metrics are reported as percentages (%).
+Moreover, we annotate UI differences using element–attribute associations, specifying the UI element being modified and the corresponding attribute change (e.g., a button with a position change), to capture real-world UI variants with concurrent modifications.
+On average, each instance contains 1.53 element-level and 1.99 attribute-level annotations, spanning 19 element types and 14 attribute types.
+We further group the element types by their primary functional roles (Figure 3(a)): Interactive Control, which directly elicits intentional user actions; Content & State Display, which primarily conveys information, semantics, or system state; and Container & Layout Structure, which organizes spatial layout and guides attention at a structural level.
+WiserUI-Bench also exhibits diversity in the interpretations used to explain user behavior, as shown in Figure 3(b), with an average of 2.28 interpretations per instance.
+The tagged UX laws are grouped into three cognitive dimensions of UX, perception, memory, and action, following Norman (2007), with four laws assigned to each.
 4 Task Description
-Proper evaluation of a model’s multimodal understanding of UI/UX design requires testing not only whether the model can identify the more effective design in guiding user behavior, but also whether it can explain why, grounding its reasoning in the visual and contextual cues.
-To this end, we define two independent tasks, decision and explanation, as illustrated in Figure 2, both essential for demonstrating genuine understanding.
-Success requires identifying key visual differences between two UIs, as well as inferring the likely behavioral goal from context (e.g., encouraging sign-ups) and how the interface supports that goal.
+Proper evaluation of a model’s multimodal understanding of UI/UX design requires assessing not only its ability to predict which design more effectively guides user behavior, but also whether it can explain why in a post-hoc manner, grounding its reasoning in visual and behavioral cues.
+To this end, we define two main tasks, selection and interpretation, as illustrated in Figure 2, both essential for demonstrating genuine understanding.
+Success requires identifying key visual differences between two UI variants, as well as inferring the underlying behavioral goal from context (e.g., promoting sign-ups) and how the interface supports that goal.
 Task 1: UI/UX Design Selection
-The first task assesses whether the model can identify which of two UI/UX designs is more effective at influencing user behavior, with the correct answer determined by large-scale A/B testing run by real companies.
+The first task assesses whether the model can identify which of two UI images is more effective at influencing user behavior, with the correct answer determined by large-scale A/B testing run by real companies.
 Although the task is framed as a binary choice, solving it effectively requires nuanced visual reasoning beyond surface-level comparison.
-Task 2: UI/UX Design Rationale Alignment
-The second task evaluates a model’s reasoning ability in depth.
+Task 2: UI/UX Design Interpretation
+The second task evaluates a model’s post-hoc interpretation ability on the observed results.
 Unlike the first task, it presents the winning UI/UX design upfront and asks the model to explain the factors contributing to its effectiveness in guiding user behavior.
-This setup mirrors the condition under which expert rationales were made, providing a fair basis for comparison.
-The model’s explanation is then compared to expert-curated key rationales, and the degree of alignment indicates the quality of its UI/UX reasoning at an expert-level.
+The model’s explanation is then compared to expert-curated key interpretations, with the degree of alignment serving as a measure of the quality of its expert-level interpretive understanding.
 5 Experiments
@@
 We evaluate a diverse set of MLLMs, spanning both proprietary and open-source models.
-On the proprietary side, we include o1 (OpenAI et al.
-2024b), GPT-4o (OpenAI et al.
-2024a), Claude 3.5 Sonnet (Anthropic 2024), all noted for their strong visual reasoning.
-For open-source models, we evaluate widely adopted models such as Qwen2.5-VL (7B, 32B) (Bai et al.
-2025), InternVL-2.5 (8B, 38B) (Chen et al.
-2025), LLaVA-NeXT-7B (Liu et al.
-2024a) and LLaVA-OneVision-7B (Li et al.
-2024b).
+On the proprietary side, we include o1 (OpenAI et al., 2024b), GPT-4o (OpenAI et al., 2024a), Claude 3.5 Sonnet (Anthropic, 2024), all noted for their strong visual reasoning.
+For open-source models, we evaluate widely adopted models such as Qwen2.5-VL (7B, 32B) (Bai et al., 2025), InternVL-2.5 (8B, 38B) (Chen et al., 2025), LLaVA-NeXT-7B (Liu et al., 2024a) and LLaVA-OneVision-7B (Li et al., 2024a).
+Refer to Appendix B.1 for details.
 5.1 UI/UX Design Selection
-Metrics
-Model accuracies are reported across input-order variations of each UI image pair, reflecting the multi-image input setting.
+5.1.1 Evaluation Metrics
+We report model accuracies across input-order variations of each UI image pair, reflecting the multi-image input setting.
 Specifically, we report: (1) First Accuracy (FA), when the more effective UI/UX design is presented first in the input; (2) Second Accuracy (SA), when it is presented second; and (3) Average Accuracy (AA), the mean of FA and SA, reflecting overall correctness.
 Finally, we compute (4) Consistent Accuracy (CA), which checks whether the model selects the same, correct UI across both input orders of each pair.
-This metric isolates the model’s core visual reasoning ability by ensuring that its choices are based on content rather than position bias, a well-known issue in pairwise comparison settings (Wang et al.
-2025b; Zhang et al.
-2023).
+This metric isolates the model’s core visual reasoning ability by ensuring that its choices are based on content rather than position bias, a well-known issue in pairwise comparison settings (Wang et al., 2025b; Zhang et al., 2023).
 (Random guessing yields a CA of 25%: since there’s a 50% chance of being correct in each order.) All evaluations are conducted over three independent runs, and we report the averaged results.
-Model FA SA AA CA
-Random 50.00 50.00 50.00 25.00
-o1 16.56 97.78 57.17 15.56
-GPT-4o 31.89 88.33 60.11 30.11
-Claude 3.5 Sonnet 26.11 86.67 56.39 24.22
-Qwen2.5-VL-32B 33.67 81.44 57.56 31.00
-Qwen2.5-VL-7B 23.00 77.89 50.44 12.89
-InternVL-2.5-38B 55.67 60.00 57.83 34.56
-InternVL-2.5-8B 58.33 45.67 52.00 21.33
-LLaVA-NeXT-7B 54.44 40.56 47.50 10.78
-LLaVA-OneVision-7B 19.78 81.56 50.67 10.44
-Table 2: Experimental results for the UI/UX design selection task on WiserUI-Bench. Best values are in bold; second-best are underlined. All metrics are reported as percentages (%).
-Results
-Table 2 summarizes model performance on the UI/UX design selection task.
-All AA scores barely clear the 50% random baseline for a binary choice.
-Moreover, they are inflated by strong position bias, especially in proprietary models and the Qwen2.5-VL family, which tend to pick the second image.
-When we examine CA, the order-invariant metric, performance drops back to near-random levels.
-Interestingly, Qwen2.5-VL-32B and InternVL-2.5-38B score slightly higher than proprietary models on CA, but still fall short of being adequate.
-Contrary to expectations, o1 shows the strongest bias and a lower CA than GPT-4o, suggesting that self-reasoning training of o1 does not transfer to this task.
-Smaller (7B/8B) models perform significantly worse than larger ones, confirming the scaling trend.
-Overall, current MLLMs struggle to consistently select the more effective UI/UX design in real-world UI pairs.
-Figure 5: Qualitative results produced by o1, GPT-4o, and Qwen2.5-VL-32B on the two core tasks.
-(Left) Qwen2.5-VL-32B and GPT-4o consistently selected the correct and incorrect variants, respectively, whereas o1 showed position bias.
-(Right) o1 successfully captured all key rationales behind the winning design.
-Green/blue highlights mark alignment with expert rationales.
-5.2 UI/UX Design Rationale Alignment
-Metrics
-To judge how well a model’s generated rationale explains the observed outcome, we compare it to a set of expert-curated key rationales per instance.
-We report two metrics: Rationale-wise accuracy, which checks model agreement with each expert rationale separately, and Instance-wise accuracy, which counts an instance as correct if the model matches any rationale.
-Alignment is assessed using a GPT-4o-based evaluator (Liu et al.
-2023), validated against human judgments on 1,000 random samples (70.8% accuracy, Cohen’s kappa 0.42).
-We also report BLEU, ROUGE, METEOR, and BERTScore (Papineni et al.
-2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al.
-2020) in Supplementary.
-Model Rationale-wise Instance-wise
-o1 64.18 78.33
-GPT-4o 50.15 66.67
-Claude 3.5 Sonnet 63.16 78.67
-Qwen2.5-VL-32B 60.23 73.00
-Qwen2.5-VL-7B 43.71 61.67
-InternVL-2.5-38B 45.32 59.33
-InternVL-2.5-8B 35.09 51.67
-LLaVA-NeXT-7B 11.99 21.67
-LLaVA-OneVision-7B 21.78 33.33
-Table 3: Results for the UI/UX design rationale alignment task on WiserUI-Bench, figures in percentages (%).
-Results
-Table 3 reports performance on the UI/UX design rationale alignment task.
-o1 and Claude 3.5 Sonnet stand out with expert-level explanations, showing strong post-hoc reasoning once the correct answer is given, even though they noticeably fell short on selection when the answer was not avaliable in Task 1.
-GPT-4o performs worse in this setting, while Qwen2.5-VL-32B remains solid and InternVL-2.5-38B drops.
-Qwen2.5-VL-7B shows relatively strong alignment, while other 7B/8B models continue to underperform.
-The consistent gap between instance-wise and rationale-wise scores indicates that while models can often align with part of the expert reasoning, they rarely capture it fully.
-6 Analyses
-6.1 Qualitative Analysis
-We analyzed differences in model outputs in depth.
-The left part of Figure 5 illustrates an example from the UI/UX design selection task.
-Variant 1, the actual A/B test winner with larger product images that are more likely to capture users’ attention, was correctly chosen by Qwen2.5-VL-32B, while other models failed to do so.
-When the correct variant appeared first in the input, Qwen2.5-VL-32B provided appropriate reasoning, demonstrating a solid understanding of UI/UX.
-In contrast, both o1 and GPT-4o incorrectly attributed their choice to an increased number of items as the reason.
-Additionally, o1 exhibited position bias.
-The right part of Figure 5 shows a case from the UI/UX design rationale alignment task.
-Given that the UI on the right was the A/B test-verified winner, experts identified the key reasons for its effectiveness as simplified background and enlargened demo button.
-both of which models must explicitly capture to provide an accurate explanation.
-o1 succeeds by directly highlighting these key points.
-GPT-4o correctly identifies the relevant UI elements, but provides an incorrect or vague explanation for their impact.
-Qwen2.5-VL-32B correctly notes the minimalist design, but stops at identifying the button, missing the key detail of its enlargement and resulting in an incomplete explanation.
-6.2 Effect of Diverse Reasoning Strategies
-For the first task, we additionally evaluate various reasoning strategies on GPT-4o and Claude 3.5 Sonnet to assess their effectiveness, including Contrastive Chain-of-Thought (CoCoT) (Zhang et al.
-2024), Self-Refine (Madaan et al.
-2023), Duty-Distinct Chain-of-Thought (DDCoT) (Zheng et al.
-2023), and Multi-Agent Debate (MAD) (Liang et al.
-2024; Du et al.
-2023).
-For MAD, we stop and moderate the debate after the first (R1) and third (R3) rounds.
-As shown in Table 4, AA scores remain similar to or below the zero-shot setting, while CA scores often surpass it, especially for MAD (R1) across models, DDCoT on GPT-4o, and CoCoT on Claude 3.5 Sonnet.
-This gain is mainly due to reduced position bias, enabling more consistent decisions regardless of input order.
-MAD (R1) performs well by generating and moderating conflicting perspectives, which suits this human preference-driven task, benefiting from multi-perspective reasoning.
-However, excessive internal debate may hurt performance, as seen in the drop with MAD (R3).
+Experiment details, including the prompts, are provided in Appendix B.2.
 Model Method FA SA AA CA
@@
 MAD (R3) 23.33 85.44 54.39 21.44
-Table 4: Results for the UI/UX design selection task across reasoning strategies on WiserUI-Bench; metrics match Section 5.1 and are averaged over three runs.
-6.3 Rationale Alignment Across UX Aspects
-We evaluated rationale-wise accuracy in the second task by each UX dimension, as truly understanding UI/UX requires capturing the full spectrum of cognitive processing—perception, memory, and action—reflected in our rationale categorization.
-As shown in Figure 6, o1 leads on Perception, indicating strong ability to reason about what users notice in a UI.
-Claude 3.5 Sonnet performs best on Action, suggesting it is more adept at predicting how users would interact with interfaces.
-Notably, model performance converge more on Memory than on the others.
-The persistent sub-70% ceiling across all dimensions underscores the need for deeper cognitive understanding of UX as a whole.
-Figure 6: Results for the UI/UX design rationale alignment task by each UX dimension on WiserUI-Bench.
-7 Discussion and Conclusion
-Toward Behavior-Aware UI Generation: Human Study on Synthetic Pairs
-As an exploratory extension of our main study, we evaluated model predictions against a human study on synthetic UI pairs crafted to elicit specific user actions.
-We position this task and its results as a stepping stone toward automating behavior-aware UI generation.
-Prior work on UI generation with large models largely targets matching natural-language or screenshot specifications (Laurençon, Tronchon, and Sanh 2024; Si et al.
-2025; Li, Zhang, and Yang 2025) or ensuring basic functionality and visual polish (Wu et al.
-2024b; Lu et al.
-2025c), rarely examining whether the produced designs actually steer user behavior.
-Because influencing actions is the core purpose of UI design, generation must also reflect that goal.
-Preference-based tuning methods such as Direct Preference Optimization (DPO) (Rafailov et al.
-2023) may offer a plausible route.
-They enable a closed-loop pipeline without direct human supervision: (1) prompt with a behavioral goal, (2) generate candidates, (3) predict which better supports the goal, and (4) feed that signal back to refine the model (Yuan et al.
-2024).
-For this to work, models must first reliably pick the more behavior-effective design not only on real-world interfaces, as in our main study, but also on synthetic pairs.
-To test this prerequisite, we conducted a human study with 55 participants on GPT-4o-generated synthetic UI pairs (code and its screenshot) based on 10 behavioral objectives.
-Participants were shown the objective and chose the variant that better supports it, yielding a preference signal focused on behavior; the majority choice serves as its proxy.
-Models, in contrast, evaluated the same pairs without access to the goal, mirroring setup in our main study.
-See Supplementary for details.
-Table 5 shows that both proprietary and open-source MLLMs, while most performing better on CA than in main study, still remain near random on AA in this synthetic evaluation, mirroring their limitations observed on WiserUI-Bench.
-Figure 7 provides a representative example.
-Going forward, consistently identifying the more effective design in such synthetic pairs will be crucial for behavior-aware UI generation, as it supplies the scalable preference signal needed for models’ closed-loop improvement.
-We encourage further work in this direction.
-Figure 7: Example from the human study.
-Given the instruction with a behavioral goal, participants selected the better synthetic variant.
-InternVL2.5-38B failed to match the majority choice due to flawed judgment.
-Model FA SA AA CA
-GPT-4o 46.67 70.00 58.33 46.67
-Claude 3.5 Sonnet 23.33 93.33 58.33 23.33
-Qwen2.5-VL-32B 43.33 66.67 55.00 40.00
-InternVL-2.5-38B 56.67 46.67 51.67 40.00
-Table 5: Results of the models on identifying the behaviorally preferred synthetic UI variant in human study. Reported figures are averaged over three trials.
-Conclusion
+Table 3: Results of the UI/UX design selection task across reasoning strategies on WiserUI-Bench, figures in percentages (%).
+5.1.2 Results
+Overall Performance
+Table 2 summarizes model performance on the UI/UX design selection task.
+Most AA scores hover only slightly above the 50% random baseline for a binary choice.
+Moreover, they are inflated by strong position bias, especially in proprietary models and the Qwen2.5-VL family, which tend to pick the second image.
+When we examine CA, the order-invariant metric, performance drops back to near-random levels.
+Interestingly, Qwen2.5-VL-32B and InternVL-2.5-38B score slightly higher than proprietary models on CA, but still fall short of being adequate.
+Contrary to expectations, o1 shows the strongest bias and a lower CA than GPT-4o, suggesting that self-reasoning of o1 does not transfer well to this task.
+Smaller (7B/8B) models perform significantly worse than larger ones, confirming the scaling trend.
+Overall, existing MLLMs struggle to consistently predict the more effective UI/UX design in real-world UI pairs.
+Model Interpretation Instance-level
+Recall Recall
+o1 64.18 78.33
+GPT-4o 50.15 66.67
+Claude 3.5 Sonnet 63.16 78.67
+Qwen2.5-VL-32B 60.23 73.00
+Qwen2.5-VL-7B 43.71 61.67
+InternVL-2.5-38B 45.32 59.33
+InternVL-2.5-8B 35.09 51.67
+LLaVA-NeXT-7B 11.99 21.67
+LLaVA-OneVision-7B 21.78 33.33
+Table 4: Results of the UI/UX design interpretation task on WiserUI-Bench, figures in percentages (%).
+Figure 5: Results of the UI/UX design interpretation task by each UX dimension on WiserUI-Bench.
+Performance Across UI Element Types
+The UI element-wise results in Table 2 show performance variations across change type categories, while preserving overall trends: GPT-4o and InternVL-2.5-38B generally achieve the highest AA and CA, respectively.
+Focusing on CA, Mixed cases, where changes span multiple functional UI element groups, perform best across most models, suggesting that aggregated multi-functional changes provide stronger behavioral cues than single-group modifications.
+Interactive Control (IC) and Content & State Display (CSD) exhibit comparable CA and rank below Mixed, indicating relatively moderate sensitivity to action-oriented (e.g., buttons) and content-level changes (e.g., images or text).
+By contrast, Container & Layout Structure (CLS) is the most challenging category, with sharply reduced CA across most models, reflecting limited sensitivity to structural layout changes, such as card or tile-level differences.
+Notably, Qwen2.5-VL-32B and InternVL-2.5-8B perform relatively better on CLS than other models, including GPT-4o.
+Figure 6: Example responses from InternVL-2.5-38B on the UI/UX design selection task, showing model outputs under different input image orders for each case.
+Blue marks correct UI change perception, while red denotes incorrect or hallucinated perception.
+Purple indicates user behavior pattern-aligned reasoning, whereas misaligned reasoning in orange.
+Figure 7: Example responses from o1, GPT-4o, and Qwen2.5-VL-32B on the UI/UX design interpretation task. Green and blue highlight points capturing the expert interpretations fully.
+Effect of Reasoning Strategies
+We additionally evaluate various reasoning strategies on GPT-4o and Claude 3.5 Sonnet to assess their effectiveness, including Contrastive Chain-of-Thought (CoCoT) (Zhang et al., 2024), Self-Refine (Madaan et al., 2023), Duty-Distinct Chain-of-Thought (DDCoT) (Zheng et al., 2023), and Multi-Agent Debate (MAD) (Liang et al., 2024; Du et al., 2023).
+For MAD, we stop and moderate the debate after the first (R1) and third (R3) rounds.
+As shown in Table 3, AA scores remain similar to or below the zero-shot setting, while CA scores often surpass it, especially for MAD (R1) across models, DDCoT on GPT-4o, and CoCoT on Claude 3.5 Sonnet.
+This gain is mainly due to reduced position bias, enabling more consistent decisions regardless of input order.
+MAD (R1) performs well by generating and moderating conflicting perspectives, which suits this human behavior pattern-driven task, benefiting from multi-perspective reasoning.
+However, excessive internal debate may hurt performance, as seen in the drop with MAD (R3).
+5.2 UI/UX Design Interpretation
+5.2.1 Evaluation Metrics
+To judge how well a model’s generated interpretation explains the observed outcome, we compare it against expert-curated key interpretations for each instance.
+Specifically, we instruct the model to produce an unconstrained free-form explanation enumerating all relevant factors, and each expert interpretation is independently evaluated in a binary manner, whether it is fully captured by the explanation or not.
+We report two metrics: Interpretation Recall, measuring the proportion of expert interpretations captured, and Instance‑level Recall, which counts an instance as correct if at least one expert interpretation is captured.
+Binary inclusion judgments are made using a GPT-4o-based evaluator (Liu et al., 2023), validated against human judgments on 1,000 random samples (83.0% accuracy, Cohen’s kappa 0.66).
+The prompts used, along with the results of BLEU, ROUGE, METEOR, and BERTScore (Papineni et al., 2002; Lin, 2004; Banerjee and Lavie, 2005; Zhang et al., 2020) are reported in the Appendix B.3.
+5.2.2 Results
+Overall Performance
+Table 4 reports performance on the UI/UX design interpretation task.
+o1 and Claude 3.5 Sonnet exhibit relatively strong post-hoc interpretation capability when the more effective variant is provided, despite having noticeably underperformed on selection in Task 1.
+By contrast, GPT-4o and InternVL-2.5-38B perform poorly under this setting, while Qwen2.5-VL-32B performs reasonably well.
+Among smaller models, Qwen2.5-VL-7B shows relatively strong alignment, whereas other 7B/8B models underperform.
+The consistent gap between Interpretation and Instance-level Recall scores indicates that while models can often recover parts of the expert interpretations, they rarely capture them in full.
+Performance Across UX Aspects
+We analyze Interpretation Recall scores by UX dimension, examining which types of UX interpretations are correctly recovered, as a true understanding of UI/UX requires capturing the full range of cognitive aspects—perception, memory, and action.
+As shown in Figure 5, o1 leads on Perception, indicating stronger ability to explain which visual cues are likely noticed by users.
+Claude 3.5 Sonnet performs best on Action, suggesting greater sensitivity to interaction-oriented factors.
+Notably, model performance converges more on Memory than on the others.
+The persistent sub-70% ceiling across all dimensions underscores the need for deeper, more holistic modeling of user cognition in UI/UX understanding and interpretation.
+6 Discussion
+Case Study
+For the UI/UX design selection task, we analyze the success and failure modes of the strong-performing model, InternVL-2.5-38B, as shown in Figure 6, focusing on cases where the model makes consistent decisions under both input image orders.
+In Case 1, the model correctly perceives the UI change as the presence of a CTA button and selects the more effective UI/UX, providing a reason that behaviorally aligns with the user behavior pattern implied by the A/B test outcome, hypothesizing that fewer CTAs better focus user attention.
+In Case 2, although the model correctly identifies differences such as pricing badges and CTA text, it fails to select the winning variant identified by the A/B test, instead offering misaligned reasoning that higher visual saliency negatively impacts user behavior by inducing commitment pressure.
+In Case 3, the model fails at perception level, hallucinating non-existent changes instead of recognizing the actual button color difference, which ultimately leads to an incorrect final decision.
+Figure 7 shows a case from the UI/UX design interpretation task.
+The UI on the right was verified by A/B test results as the more effective design, with experts attributing its effectiveness to a simplified background and an enlarged demo button, along with their associated effects, factors that models must explicitly capture for an accurate interpretation.
+o1 succeeds by directly highlighting these key points.
+GPT-4o correctly identifies the relevant UI elements, but provides an incorrect or vague explanation for their impact.
+Qwen2.5-VL-32B correctly notes the minimalist design, but stops at identifying the button, missing the key detail of its enlargement and resulting in an incomplete explanation.
+7 Conclusion
 We introduce a new task that evaluates multimodal understanding of how design drives user actions in the critical field of UI/UX.
-To support this, we present WiserUI-Bench, a behavior-focused benchmark uniquely built from diverse, real-world A/B-tested UI variants and expert-curated key rationales, providing a reliable basis for assessment.
-Our two tasks, selection and explanation, go beyond surface-level recognition, requiring models to reason about the behavioral implications of visual design, a cognitively challenging problem.
-Extensive experiments across both current proprietary and open-source MLLMs reveal they exhibit only shallow levels of nuanced visual reasoning.
-Achieving stronger performance on this significant task may, for example, require further training on vast amounts of behaviorally grounded UI/UX data that reflect real-world dynamics.
-We envision WiserUI-Bench and our work as a step toward multimodal systems capable of not only understanding but also generating behavior-aware interfaces, aligning with human cognition and preferences to optimize user experience.
-More broadly, we hope to catalyze future research toward deeper UI/UX and human-behavior literacy in AI.
-A Benchmark for Multimodal UI/UX Design Understanding
-Supplementary Material
-Appendix A Overview
-This supplementary material includes the following sections: (B) Limitations of our work, (C) Ethics Statements for UI/UX experts and human study participants, Detailed descriptions of (D) WiserUI-Bench, (E) experiments, and (F) human study, and concludes with (G) icon attribution.
-Appendix B Limitation
-Given that user experience patterns vary across cultural contexts and societal norms, cultural biases inevitably exist in our benchmark (Karreman, Romeo, and Li 2016).
-However, accounting for such diversity is also part of the benchmark’s intent, making it a more realistic and challenging evaluation setting.
+To support this, we present WiserUI-Bench, a behavior-focused benchmark uniquely built from diverse, real-world A/B‑tested UI variants and expert-curated key interpretations, providing a reliable basis for assessment.
+Our two tasks, selection and interpretation, go beyond surface-level recognition, requiring models to reason about the behavioral implications of visual design, a cognitively challenging problem.
+Extensive experiments across both proprietary and open-source MLLMs reveal they exhibit only shallow levels of behavior-aware visual reasoning.
+Achieving stronger performance on this significant task, particularly across the analyzed UI elements or UX dimensions, may require further training on large-scale, behaviorally grounded UI/UX data that reflects real-world user dynamics.
+We envision WiserUI-Bench and our work as a step toward multimodal systems capable of not only understanding but also generating visual interfaces grounded in user behavior contexts, aligning with human cognition to optimize user experience.
+More broadly, we hope to catalyze future research toward deeper UI/UX and human behavior literacy in AI.
+8 Limitations
+Given that user experience patterns vary across cultural contexts and societal norms, cultural biases inevitably exist in our benchmark (Karreman et al., 2016).
 Also, interactive UIs were not extensively explored in our work.
 Although the number of 300 UI image pairs may seem limited, publicly available real-world A/B test results are extremely rare, as companies rarely disclose them.
-Despite the dataset size, it covers a wide range of diverse UI changes, each accompanied by multiple expert-validated rationales.
-Appendix C Ethics Statement
-To ensure expert-level annotation quality, we recruited three annotators with professional UI/UX backgrounds, all currently employed at global IT companies specializing UX and AI.
-They were clearly informed of the academic nature of the task and voluntarily participated in the study.
-Each annotator received monetary compensation equivalent to approximately \$225 USD per person, which reflects fair market payment for domain-specific labor.
-Payments were issued within 48 hours of annotation completion.
-All annotators were instructed to focus solely on task-relevant elements and were not exposed to or asked to label any sensitive, offensive, or triggering content.
-The human study was also conducted with strict ethical considerations.
-To minimize biases, we employed a double-blind setup.
-All 55 participants voluntarily took part in the survey, fully informed that the study was for academic purposes.
-Additionally, the survey contained no explicit or problematic content, ensuring a safe and respectful research environment.
-Appendix D WiserUI-Bench Details
-D.1 Image Preprocess
-During benchmark construction process, we observed that some UI images from our data sources included manual indicators, such as circled regions, numbered markers, and arrows, highlighting specific UI elements related to reasoning points from A/B test results (see Figure J).
+Moreover, independently conducting production-level UI experiments at a comparable level and scale was practically infeasible.
+However, as the first benchmark in this area, we focused on reliability, diversity, and annotation quality, covering a wide range of real-world UI changes and providing expert-curated key interpretations.
+9 Ethical Considerations
+To ensure expert-level annotation quality, we recruited three annotators with professional UI/UX backgrounds who are currently employed at global IT companies specializing in UX and AI.
+We clearly informed them of the academic nature of the task, and they voluntarily participated in the study.
+We provided each annotator with monetary compensation equivalent to approximately \$225 USD, reflecting fair market rates for domain-specific labor, and issued payments within 48 hours of annotation completion.
+We instructed annotators to focus solely on task-relevant elements and ensured that they were not exposed to or asked to label any sensitive, offensive, or triggering content.
+Our benchmark does not include any private, user-identifiable, or confidential information.
+All materials were collected from publicly released sources originally disclosed by the originating platforms or companies.
+Nevertheless, to avoid downstream commercial use beyond the original disclosure scope, we plan to release the benchmark under a non-commercial research license.
+Appendix A WiserUI-Bench Details
+A.1 Curation Details
+Image Preprocess
+During benchmark construction process, we observed that some UI images from our data sources included manual indicators, such as circled regions, numbered markers, and arrows, highlighting specific UI elements related to reasoning points from A/B test results (see Figure 17).
 While helpful for human interpretation, these indicators introduced artificial attention cues that could potentially bias model evaluation.
@@
 This process ensured a cleaner, more naturalistic dataset for fair visual reasoning evaluation.
-D.2 Rationale Annotation
-As shown in Figure L, three UI/UX experts are presented with a pair of UI images, one of which was designated as the winner based on A/B test results, along with a corresponding source website link.
-Experts were instructed to write key rationales explaining why the winning version supported user actions more effectively than the alternative.
+Interpretation Annotation
+As shown in Figure 20, three UI/UX experts are presented with a pair of UI images, one designated as the winner based on A/B test results, along with a corresponding source website link.
+Experts were instructed to provide key interpretations explaining why the winning version supported user actions more effectively than the alternative.
 We engaged experts for annotation, as user behavior is often subconscious and shaped by implicit context, making evaluation by non-experts inherently unreliable.
-These rationales underwent multiple rounds of review to ensure that no important reason was omitted and no irrelevant reason was included.
-Expert annotators were guided by the original sources’ textual description which often highlights key design intentions, and two types of theoretical heuristics selected for their robustness and relevance to user behavior analysis: (1) the Nielsen Norman 10 Usability Heuristics (Nielsen 1994), serving as broad conceptual criteria, and (2) 12 curated UX laws that are still widely applied in practice.
+To distill salient behavioral drivers, we retained only interpretations independently identified by at least two experts, resulting in 684 final interpretations (562 with unanimous agreement and 122 with two-expert agreement) from an initial pool of 845 candidates.
+Expert annotators were initially guided by the original sources’ textual description which often highlights key design intentions, as well as by a robust theoretical heuristic widely used in usability research, the Nielsen Norman 10 Usability Heuristics (Nielsen, 1994), which served as broad conceptual criteria.
+The resulting interpretations were then consolidated and mapped to 12 curated UX laws that remain widely applied in practice.
 These UX laws provide concrete, actionable guidance across diverse aspects of user experience, which can be categorized into three cognitive dimensions: Perception, Memory, and Action.
@@
 Emotional design focuses on creating products that evoke emotions and foster positive user experience by engaging users at three cognitive levels: visceral, behavioral, and reflective.
-The full list of UX laws and their corresponding classifications is presented in Table A.
-D.3 Statistics
-Figure K presents the distribution of WiserUI-Bench across page types, industry domains, device types, and types of UI element changes.
+The full list of UX laws and their corresponding classifications is presented in Table 5.
+A.2 Statistics Details
+Figure 18 presents the distribution of WiserUI-Bench across page types, industry domains, device types.
 For clarification on the page type classification, the homepage refers to the main entry point of a website, typically offering broad navigation and general content.
 In contrast, a landing page is usually a standalone page optimized for a specific user action, such as signing up or making a purchase.
-For clarification on the types of UI element changes, detailed definitions are provided in Table B.
-Appendix E Experiment Details
-E.1 Setup
-Model Configuration
+Full distribution of UI change attribute types is shown in Figure 19.
+For clarification, element-level counts reflect the number of distinct UI components affected, while attribute-level counts capture the total number of modified attributes across all components.
+For example, an instance annotated as <Input Field: Padding, Color / Button: Color> is counted as 2 element changes and 3 attribute changes.
+Appendix B Experiment Details
+B.1 Setup
+B.1.1 Model Configuration
 For proprietary MLLMs, we accessed the models via their official API endpoints as follows:
-  • o1 (OpenAI et al. 2024b):
+  • o1 OpenAI et al. (2024b):
     o1-2024-12-17
-  • GPT-4o (OpenAI et al. 2024a):
+  • GPT-4o OpenAI et al. (2024a):
     gpt-4o-2024-08-06
-  • Claude 3.5 Sonnet (Anthropic 2024):
+  • Claude 3.5 Sonnet Anthropic (2024):
     claude-3-5-sonnet-20240620
 For open-source models, we used publicly available weights from the Huggingface Hub [^3] https://huggingface.co/models. Specific model versions are as follows:
-  • Qwen2.5-VL-32B (Bai et al. 2025):
+  • Qwen2.5-VL-32B Bai et al. (2025):
     Qwen/Qwen2.5-VL-32B-Instruct
-  • Qwen2.5-VL-7B (Bai et al. 2025):
+  • Qwen2.5-VL-7B Bai et al. (2025):
     Qwen/Qwen2.5-VL-7B-Instruct
-  • InternVL-2.5-38B (Chen et al. 2025):
+  • InternVL-2.5-38B Chen et al. (2025):
     OpenGVLab/InternVL2_5-38B
-  • InternVL-2.5-8B (Chen et al. 2025):
+  • InternVL-2.5-8B Chen et al. (2025):
     OpenGVLab/InternVL2_5-8B
-  • LLaVA-NeXT-7B (Liu et al. 2024a):
+  • LLaVA-NeXT-7B Liu et al. (2024a):
     llava-hf/llava-v1.6-mistral-7b-hf
-  • LLaVA-OneVision-7B (Li et al. 2024b):
+  • LLaVA-OneVision-7B Li et al. (2024a):
     llava-hf/llava-onevision-qwen2-7b-ov-hf
 The temperature for all models was set to 0.2.
-Resources
+B.1.2 Resources
 All experiments using open-source models were conducted on a Linux machine (Ubuntu 20.04.6 LTS) equipped with 4 NVIDIA RTX A6000 GPUs, each with 48GB VRAM.
 We used the following software libraries and frameworks: transformers version 4.53.2 and vllm version 0.9.2.
-E.2 UI/UX Design Selection
-Prompts
-Figures A through G show the prompts used, including those designed for diverse reasoning strategies.
-Token Counts
-Token counts across reasoning strategies on proprietary models are presented in Table C.
-Standard Deviation
-Table D presents the standard deviation across three experiment runs for each model and reasoning method.
+B.2 UI/UX Design Selection
+B.2.1 Prompts
+Figures 8 through 14 show the prompts used, including those designed for diverse reasoning strategies.
+B.2.2 Token Counts
+Token counts across reasoning strategies on proprietary models are presented in Table 6.
+B.2.3 Standard Deviation
+Table 7 presents the standard deviation across three experiment runs for each model and reasoning method.
 For metric abbreviation, FS and SS denote the standard deviation when the first or second UI image input was the more effective one, while AS and CS refer to the standard deviation of total correct and both correct cases, respectively.
-Additional Qualitative Results
-Figure N and O present the additional qualitative examples across different models. Figure P shows an example comparing various reasoning strategies executed with GPT-4o.
-E.3 UI/UX Design Rationale Alignment
-Prompts
-Figure H presents the prompt used for models’ rationale generation, and Figure I presents the prompt used for its GPT-4o-based evaluation.
-Results on Text Similarity Metrics
-Alongside rationale-wise and instance-wise accuracies from GPT-4o-based evaluation, we report BLEU, ROUGE, METEOR, and BERTScore as well.
-(Papineni et al.
-2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al.
-2020) in Table E.
-Additional Qualitative Results
-Figure Q presents an additional qualitative example across different models.
-Appendix F Human Study Details
-Figure M displays the interface used for human evaluation, where participants select which synthetic UI better adheres to the given instructions, including the specified behavioral goal.
-Table F presents the results of the human study, showing the distribution of participant responses across 10 A/B questions.
-For each question, the option chosen by the majority is treated as a proxy for human preference and is highlighted in bold.
-Results of standard deviation across three experiment runs are shown in Table G.
-Figure R provides an additional qualitative example of how models performed in identifying the preferred UI variant.
-Appendix G Icon Attribution
+B.2.4 Additional Case Study
+Figure 21 and 22 present the additional case studies across different models. Figure 23 shows an example comparing various reasoning strategies executed with GPT-4o.
+B.3 UI/UX Design Interpretation
+B.3.1 Prompts
+Figure 15 presents the prompt used for models’ interpretation generation, and Figure 16 presents the prompt used for its GPT-4o-based evaluation.
+B.3.2 Results on Text Similarity Metrics
+Alongside Interpretation Recall and Instance-level Recall from GPT-4o-based evaluation, we report BLEU, ROUGE, METEOR, and BERTScore as well.
+Papineni et al.
+(2002); Lin (2004); Banerjee and Lavie (2005); Zhang et al.
+(2020) in Table 8.
+B.3.3 Additional Case Study
+Figure 24 presents an additional case study across different models.
+Appendix C Icon Attribution
 Icons used in this paper are from Flaticon (https://www.flaticon.com), and they are attributed to the respective authors as required by Flaticon’s license.
 UX Dimension UX Law Description
-Perception Law of Common Region (Koffka 1922) Elements sharing a boundary or background are perceived as a unified group.
+Perception Law of Common Region Koffka (1922) Elements sharing a boundary or background are perceived as a unified group.
 This can be a visible border, a colored box, or a distinct region in the layout.
-Law of Proximity (Koffka 1922) Items positioned close together are seen as belonging to the same group. This mental shortcut helps users quickly identify relationships between adjacent elements.
-Law of Prägnanz (Koffka 1922) Humans naturally perceive and interpret complex images in their simplest forms. It reduces cognitive effort by helping us group and understand elements more quickly.
-Law of Similarity (Koffka 1922) People group together elements that look alike, assuming they serve a similar purpose.
+Law of Proximity (Koffka, 1922) Items positioned close together are seen as belonging to the same group. This mental shortcut helps users quickly identify relationships between adjacent elements.
+Law of Prägnanz (Koffka, 1922) Humans naturally perceive and interpret complex images in their simplest forms. It reduces cognitive effort by helping us group and understand elements more quickly.
+Law of Similarity (Koffka, 1922) People group together elements that look alike, assuming they serve a similar purpose.
 Consistency in color, shape, or size leads the user to see these items as related.
-Memory Von Restorff Effect (Von Restorff 1933) People remember a distinct element more easily than others that blend in.
+Memory Von Restorff Effect (Von Restorff, 1933) People remember a distinct element more easily than others that blend in.
 This is also known as the Isolation Effect, emphasizing our attention on anything unusual.
-Serial Position Effect (Ebbinghaus, Ruger, and Bussenius 1913) Users tend to remember the first and last items in a series more than those in the middle.
+Serial Position Effect (Ebbinghaus et al., 1913) Users tend to remember the first and last items in a series more than those in the middle.
 This influences how they recall sequential information or actions.
-Miller’s Law (Miller 1956) On average, people can hold only about seven items in their working memory at once. Going beyond this limit causes cognitive overload.
-Jakob’s Law (Nielsen 2000) Users expect new interfaces to operate similarly to those they’ve used before. Leveraging established patterns lowers cognitive load and accelerates adoption.
-Action Hick’s Law (Hick 1952) Decision time rises with the number and complexity of choices. Too many options can overwhelm users and slow their actions.
-Pareto Principle (Pareto 1919) Roughly 80% of outcomes stem from 20% of causes. Focusing on the most impactful features or tasks delivers the greatest improvement.
-Fitts’s Law (Fitts and Peterson 1964) The time required to reach a target depends on its size and distance. Larger, closer elements are easier to access quickly and accurately.
-Goal-Gradient Effect (Hull 1932) Motivation increases as people get closer to finishing a task. Providing clear progress markers encourages users to continue and complete it faster.
-Table A: Overview of UX laws used for rationale annotation, and their categorization by cognitive dimension (Perception, Memory, Action).
-Type Definition
-Content & Copy All textual elements that convey meaning or guide the user, including headlines, body text, labels, error messages, promotional language, and social proof content.
-Visual & Style Visual design elements such as color schemes, typography, spacing, icon sizes, and overall aesthetic consistency that shape the look and feel of the interface.
-Controls & Inputs Interactive UI components that allow users to perform actions, including buttons, toggles, dropdowns, sliders, search fields, and other input elements.
-Layout & Structure The spatial organization of UI components, covering alignment, grouping, column or grid usage, visual hierarchy, and structural consistency across screens.
-Navigation & Flow Interface elements that support user movement and orientation within the product, such as menus, sidebars, breadcrumbs, tabs, and filter panels.
-Media & Visual Assets Non-textual content like images, icons, thumbnails, banners, and videos used to support communication, branding, or user engagement.
-Pricing & Promotion How pricing details, discounts, promotional messages, limited-time offers, or benefits like free trials are visually and textually communicated to the user.
-Trust & Social Proof UI elements that aim to build user trust and credibility, including customer reviews, ratings, trust badges, security seals, guarantees, and testimonials.
-Forms & Form UX Design and usability aspects of data entry forms, such as number and arrangement of fields, required indicators, input validation, and overall ease of completion.
-Status & Feedback Visual and textual responses to user actions or system states, including loading indicators, progress bars, tooltips, error messages, confirmations, and alerts.
-Table B: Definitions of UI element change types.
+Miller’s Law (Miller, 1956) On average, people can hold only about seven items in their working memory at once. Going beyond this limit causes cognitive overload.
+Jakob’s Law (Nielsen, 2000) Users expect new interfaces to operate similarly to those they’ve used before. Leveraging established patterns lowers cognitive load and accelerates adoption.
+Action Hick’s Law (Hick, 1952) Decision time rises with the number and complexity of choices. Too many options can overwhelm users and slow their actions.
+Pareto Principle (Pareto, 1919) Roughly 80% of outcomes stem from 20% of causes. Focusing on the most impactful features or tasks delivers the greatest improvement.
+Fitts’s Law (Fitts and Peterson, 1964) The time required to reach a target depends on its size and distance. Larger, closer elements are easier to access quickly and accurately.
+Goal-Gradient Effect (Hull, 1932) Motivation increases as people get closer to finishing a task. Providing clear progress markers encourages users to continue and complete it faster.
+Table 5: Overview of UX laws used for interpretation annotation, and their categorization by cognitive dimension (Perception, Memory, Action).
 Models Methods Input Tokens Output Tokens
@@
 MAD (R3) 13853.4 ($\pm$ 7512.63) 1936.5 ($\pm$ 276.14)
-Table C: Overview of Token Counts ($\mu\pm\sigma$) in the UI/UX design selection task on WiserUI-Bench.
-Model Method FS SS AS CS
+Table 6: Overview of Token Counts ($\mu\pm\sigma$) in the UI/UX design selection task on WiserUI-Bench.
+Model Method FS $S$ AS CS
 o1 Zero-Shot 1.528 2.082 3.606 0.577
@@
 LLaVA-OneVision-7B Zero-Shot 4.041 2.309 4.359 3.512
-Table D: Results on standard deviation across three experiment runs in the UI/UX design selection task for each model and reasoning strategy on WiserUI-Bench.
+Table 7: Results on standard deviation across three experiment runs in the UI/UX design selection task for each model and reasoning strategy on WiserUI-Bench.
 Model BLEU-2 BLEU-4 ROUGE-L METEOR BERTScore
@@
 LLaVA-OneVision-7B 0.0157 0.0043 0.0718 0.1416 0.8255
-Table E: Results of text similarity metrics per expert rationale for the UI/UX design rationale alignment task on WiserUI-Bench.
-Choice Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q10
-A 42 9 23 14 25 40 14 27 35 18
-B 13 46 32 41 30 15 41 28 20 37
-Majority Choice A B B B B A B B A B
-Table F: Distribution of responses from human participants indicating their preferred UI variant across 10 A/B test questions.
-Model FS SS AS CS
-GPT-4o 0.577 0.000 0.577 0.577
-Claude 3.5 Sonnet 0.577 0.577 0.577 0.577
-Qwen2.5-VL-32B 0.577 0.577 1.000 0.000
-InternVL-2.5-38B 0.577 0.577 0.577 1.000
-Table G: Results on standard deviation across three experiment runs on identifying the behaviorally preferred synthetic UI variant in human study.
+Table 8: Results of text similarity metrics per expert interpretation for the UI/UX design interpretation task on WiserUI-Bench.
 You are an expert in designing UI/UX for web/apps.
@@
 You should end your answer with following the format (No bold, etc): More effective: <First/Second>
-Figure A: Prompt used for the UI/UX design selection task with zero-shot strategy.
+Figure 8: Prompt used for the UI/UX design selection task with zero-shot strategy.
 You are an expert in designing UI/UX for web/apps.
@@
 You should end your answer with the following format (No bold, etc): More effective: <First/Second>
-Figure B: Prompt used for the UI/UX design selection task with CoCoT strategy.
+Figure 9: Prompt used for the UI/UX design selection task with CoCoT strategy.
 You are an expert in designing UI/UX for web/apps.
@@
 Previous answer: {previous_answer}
-Figure C: Prompt used for the UI/UX design selection task with Self-Refine strategy on the reviewing step. Prompt inputs are in bold.
+Figure 10: Prompt used for the UI/UX design selection task with Self-Refine strategy on the reviewing step. Prompt inputs are in bold.
 You are an expert in designing UI/UX for web/apps.
@@
 You should end your answer with following the format (No bold, etc): More effective: <First/Second>
-Figure D: Prompt used for the UI/UX design selection task with Self-Refine strategy on the improving step. Prompt inputs are in bold.
+Figure 11: Prompt used for the UI/UX design selection task with Self-Refine strategy on the improving step. Prompt inputs are in bold.
 You are an expert in designing UI/UX for web/apps.
@@
 You should end your answer with the following format (No bold, etc): More effective: <First/Second>
-Figure E: Prompt used for the UI/UX design selection task with DDCoT strategy.
+Figure 12: Prompt used for the UI/UX design selection task with DDCoT strategy.
 You are an expert in designing UI/UX for web/apps.
@@
 The opponent proposed that: {opponent_opinion}
-Figure F: Prompt used for the UI/UX design selection task with MAD strategy on the debate step arguing the first/second version better drives user actions. Prompt inputs are in bold.
+Figure 13: Prompt used for the UI/UX design selection task with MAD strategy on the debate step arguing the first/second version better drives user actions. Prompt inputs are in bold.
 You are an expert in designing UI/UX for web/apps.
@@
 You should end your answer with following the format (No bold, etc): More effective: <First/Second>
-Figure G: Prompt used for the UI/UX design selection task with MAD strategy on the moderation step. Prompt inputs are in bold.
+Figure 14: Prompt used for the UI/UX design selection task with MAD strategy on the moderation step. Prompt inputs are in bold.
 You are an expert in designing UI/UX for web/apps.
 The two screenshots show two different versions of the same page.
-If the SECOND version has proven to be more effective UI/UX design that leads to better user experience and conversion, provide all possible and appropriate for its superiority.
-Figure H: Prompt used for the UI/UX design rationale alignment task.
-Given two statements are reasons why one UI is better than another in terms of guiding user behavior effectively. Statement 1: {model_reason}
+If the SECOND version has proven to be more effective UI/UX design that leads to better user experience and conversion, provide all possible and appropriate interpretations for its superiority.
+Figure 15: Prompt used for the UI/UX design interpretation task.
+Given two statements are interpretations why one UI is better than another in terms of guiding user behavior effectively. Statement 1: {model_reason}
 Statement 2: {reference_reason} Your task is to assess whether the first statement contains the same reasoning and essential content as the second statement.
@@
 Otherwise, answer with ’No’.
-Figure I: Prompt used for GPT-4o-based evaluation on the UI/UX design rationale alignment task. Prompt inputs are in bold.
-Figure J: Example of visual indicator removal in WiserUI-Bench.
+Figure 16: Prompt used for GPT-4o-based evaluation on the UI/UX design interpretation task. Prompt inputs are in bold.
+Figure 17: Example of visual indicator removal in WiserUI-Bench.
 Left: Original image with markers added by data source website.
 Right: After inpainting, with markers removed while preserving UI elements.
-Figure K: Detailed statistics of WiserUI-Bench on (a) page type, (b) industry domain, (c) device type, (d) UI element change type
-Figure L: A screenshot of the rationale annotation interface for UI/UX experts based on original data sources and guidelines.
-While writing the key rationales, they also selected the corresponding UX law.
-Figure M: A screenshot of the human study interface. Each participant selects the UI variant they believe better supports the given instruction, including the specified behavioral goal.
-Figure N: A qualitative example from the UI/UX design selection task on WiserUI-Bench: Both Claude 3.5 Sonnet and InternVL-2.5-38B selected the A/B test winner.
-Figure O: A qualitative example from the UI/UX design selection task on WiserUI-Bench: GPT-4o selected the A/B test winner, whereas LLaVA-NeXT-7B did not.
-Figure P: A qualitative example from the UI/UX design selection task on WiserUI-Bench, showcasing the effect of diverse reasoning strategies executed on GPT-4o.
-Figure Q: A qualitative example from the UI/UX design rationale alignment task on WiserUI-Bench: InternVL-2.5-8B captured only the memory rationale, while Qwen2.5-VL-32B identified two key rationales.
-The highlights in the models’ responses indicate the parts that align with the expert-curated key rationales.
-Figure R: A qualitative example from the UI/UX design selection on human study: InternVL2.5-38B failed to match the majority choice, whereas Claude 3.5 Sonnet succeeded.
-For fair evaluation, we inserted the same images into the placeholder area of each generated UI, as shown above.
+Figure 18: Detailed statistics of WiserUI-Bench on (a) page type, (b) industry domain, (c) device type
+Figure 19: Detailed statistics of WiserUI-Bench on UI change attribute type
+Figure 20: A screenshot of the interpretation annotation interface for UI/UX experts based on original data sources and guidelines.
+After writing the key interpretations, they also selected the most relevant UX law.
+Figure 21: A qualitative example from the UI/UX design selection task on WiserUI-Bench: Both Claude 3.5 Sonnet and InternVL-2.5-38B selected the A/B test winner.
+Figure 22: A qualitative example from the UI/UX design selection task on WiserUI-Bench: GPT-4o selected the A/B test winner, whereas LLaVA-NeXT-7B did not.
+Figure 23: A qualitative example from the UI/UX design selection task on WiserUI-Bench, showcasing the effect of diverse reasoning strategies executed on GPT-4o.
+Figure 24: A qualitative example from the UI/UX design interpretation task on WiserUI-Bench: InternVL-2.5-8B captured only the memory part, while Qwen2.5-VL-32B identified two key interpretations.
+The highlights in the models’ responses indicate the parts that align with the expert-curated key interpretations.